FLOSS01: Logbook

FLOSS01: Chronological Logbook Entries


Return to Logbook Top Page
Entry Date Title Site Author #Graphics
180 Wed 03-Apr-2002Teardowntowermaclean
179 Wed 03-Apr-2002Tower soil samplestowermaclean
178 Wed 27-Mar-2002Radiometers and soil sensors removed from all sitesall stnsdelany
177 Wed 27-Mar-2002Soil coretowerdelany
176 Mon 18-Mar-2002Kryptons bad at PAM sutessemmer
175 Sun 17-Mar-200220m prop down againtowermaclean
174 Thu 14-Mar-2002Daily status, 3/14/02semmer
173 Wed 13-Mar-2002rad & kryptons cleanedall stnsoncley
172 Wed 13-Mar-2002rad cleanedtoweroncley
171 Wed 13-Mar-2002soil samplestower
170 Wed 13-Mar-2002cal gasestoweroncley
169 Wed 13-Mar-2002krytpons cleanedtoweroncley
168 Wed 13-Mar-2002several thingstoweroncley
167 Tue 12-Mar-2002Daily status, 3/12/02semmer
166 Sat 09-Mar-2002cosmos analog channels flatlinedtowermaclean
165 Fri 08-Mar-200220 meter prop outsemmer
164 Mon 04-Mar-20024-component aspiration fans disconnected 3/2stn3militzer
163 Mon 04-Mar-2002Sta 3 Outages 3/2, 3/3stn3militzer
162 Fri 01-Mar-2002Soil core at towertowerdelany
161 Fri 01-Mar-2002Replace Kr at Sagestn3delany
160 Wed 27-Feb-2002Guy tension/chem sheltertowerdelany
159 Wed 27-Feb-200220 m propvane reinstalledtowerdelany
158 Wed 27-Feb-2002Rain guage inspectiontowerdelany
157 Wed 27-Feb-2002Soil cores at Sage and Baredelany
156 Tue 26-Feb-2002Removed Licor 7500towerdelany
155 Mon 25-Feb-2002Reinstalled 1 m propvanetowerdelany
154 Fri 22-Feb-2002Axis webcam settingstowermaclean
153 Fri 22-Feb-2002Kr and Rad cleanall stnsdelany
152 Fri 22-Feb-2002Remove 1m propvanetowerdelany
151 Thu 21-Feb-2002Guy wire tension adjustmenttowerdelany
150 Thu 21-Feb-2002Ultrasonic snow depth checktowerdelany
149 Thu 21-Feb-2002Cal gas pressure log for #5 & #6towerdelany
148 Thu 21-Feb-2002Radiometer visitsall stnsdelany
147 Wed 20-Feb-2002Soil moisture profiletowerdelany
146 Wed 20-Feb-2002Soil samples at sage, bare and towerall stnsdelany
145 Wed 20-Feb-2002Change cylinders, check flow ratestowerdelany
144 Tue 19-Feb-2002towerdelany
143 Tue 19-Feb-2002new kryptonstn2oncley
142 Tue 19-Feb-2002new kryptonstn3oncley
141 Mon 18-Feb-2002soil samplestoweroncley
140 Mon 18-Feb-2002thermocouples installedtoweroncley
139 Mon 18-Feb-2002kryptons cleanedall stnsoncley
138 Mon 18-Feb-2002rads cleanedall stnsoncley
137 Mon 18-Feb-2002power system inventoryall stnsoncley
136 Mon 18-Feb-2002low? CO2 cylinder changedtoweroncley
135 Sun 17-Feb-200220m prop speed decoder outtowermaclean
134 Sun 17-Feb-2002Lost webcam pictures Feb14-16towermaclean
133 Tue 12-Feb-2002all stnsdelany
132 Sat 09-Feb-2002Weekly Status, 2-9 Feb 2002militzer
131 Wed 06-Feb-2002network cameratowermaclean
130 Fri 25-Jan-2002Daily status, 1/25/02semmer
129 Thu 24-Jan-2002Daily status, 1/24/02semmer
128 Wed 23-Jan-2002Daily status, 1/23/02semmer
127 Wed 23-Jan-2002Daily status, 1/22/02semmer
126 Tue 22-Jan-2002Daily status, 1/18/02horst
125 Tue 22-Jan-2002Daily status, 1/17/02horst
124 Fri 18-Jan-2002RAD domes cleaned at tower and sagesemmer
123 Fri 18-Jan-2002CAL GAS readingssemmer
122 Fri 18-Jan-2002trime at towersemmer
121 Fri 18-Jan-2002Installed web camtowermaclean
120 Fri 18-Jan-2002Added h2oV.barestn2maclean
119 Thu 17-Jan-2002Trime at sagestn3semmer
118 Thu 17-Jan-2002soil samplesall stnssemmer
117 Thu 17-Jan-2002added tubing to cal gas linestowersemmer
116 Thu 17-Jan-2002soil sample from Tony's visitall stnssemmer
115 Wed 16-Jan-2002Wrong cal used for krypton at sage sitestn3maclean
114 Wed 16-Jan-2002Daily status, 1/16/02horst
113 Tue 15-Jan-2002Daily status, 1/15/02horst
112 Mon 14-Jan-2002Daily Status, 1/14/02horst
111 Mon 07-Jan-2002Visit 01/04/2002 Task CheckList Reviewall stnsmilitzer
110 Fri 04-Jan-2002PAM Kr cleanedstn3delany
109 Fri 04-Jan-2002guy wire tension updatetowermilitzer
108 Fri 04-Jan-2002Swapped Base Disk Drivebasemilitzer
107 Fri 04-Jan-2002Snow Depth Test: Thermocouple exposed to airtowermilitzer
106 Fri 04-Jan-2002Radiometer Domes cleanedall stnsmilitzer
105 Fri 04-Jan-2002Kryptons Cleanedtowermilitzer
104 Fri 04-Jan-2002New Thermocouples Installed.towermilitzer
103 Fri 04-Jan-200220m Prop Replacedtowermilitzer
102 Thu 03-Jan-2002li7500 calibrationstoweroncley1
101 Thu 27-Dec-2001New temp variable from snow depth gaugetowermaclean
100 Wed 26-Dec-2001temperature event?all stnsoncley
99 Wed 26-Dec-2001Rsw.in(main) oddtoweroncley
98 Wed 26-Dec-2001temp sensitivitiesall stnsoncley
97 Wed 26-Dec-2001temp sensitivitytoweroncley
96 Wed 19-Dec-2001archive size and disk capacityall stnsmaclean
95 Tue 18-Dec-2001TRH fan replacedstn2militzer
94 Tue 18-Dec-2001Sage site sonic boom anglestn3militzer
93 Tue 18-Dec-2001guy wire tensiontoweroncley
92 Mon 17-Dec-2001TRH fans replacedtowermilitzer
91 Mon 17-Dec-20014-Component Radiometer Stn2 moved to Stn3stn2militzer
90 Mon 17-Dec-2001PAM Stn3 / sage site setupstn3militzer
89 Thu 13-Dec-2001calibration files for CR10Xsall stnssemmer
88 Wed 12-Dec-2001Changed Icharge monitor to Iwind/Isolarstn2militzer
87 Mon 10-Dec-2001aster rebooted several timesbaseoncley
86 Mon 10-Dec-2001Change of sw.out at maintowerdelany
85 Mon 10-Dec-2001new licor7500 convertertoweroncley
84 Mon 10-Dec-2001restarting serial_ingestbasemaclean
83 Mon 10-Dec-2001level sensors temperature errortoweroncley
82 Sun 09-Dec-2001rebooted asterall stnsoncley
81 Sun 09-Dec-2001rads cleaned?stn2oncley
80 Sun 09-Dec-2001krypton cleanedstn2oncley
79 Sun 09-Dec-2001visitorsall stnsoncley
78 Sun 09-Dec-2001rads servicedtoweroncley
77 Sun 09-Dec-2001kryptons cleanedtoweroncley
76 Sat 08-Dec-2001shallow inversion todayall stnsoncley
75 Sat 08-Dec-2001replaced last thermocouplestoweroncley
74 Sat 08-Dec-2001staked rad cablestn2oncley
73 Sat 08-Dec-2001li7500 uptoweroncley
72 Sat 08-Dec-2001TRH 703all stnsoncley
71 Sat 08-Dec-2001soil samplestoweroncley
70 Sat 08-Dec-2001serial numbers and heightsall stnsoncley
69 Sat 08-Dec-2001T/RH swapped ~11/30toweroncley
68 Sat 08-Dec-2001soil samplestoweroncley
67 Sat 08-Dec-2001gps died againbaseoncley
66 Fri 07-Dec-2001sonic interruptionsstn2oncley
65 Thu 06-Dec-2001flow rates checkedtoweroncley
64 Thu 06-Dec-2001Lower thermocouples: changes to cr21xtowermaclean
63 Thu 06-Dec-2001soil readingstoweroncley1
62 Thu 06-Dec-2001soil samplesstn2oncley
61 Thu 06-Dec-2001soil samplingtowerdelany
60 Thu 06-Dec-2001CO2 installationtowerdelany
59 Thu 06-Dec-2001radiometers and soil sensorsall stnsdelany1
58 Thu 06-Dec-2001cal gasestowerdelany1
57 Thu 06-Dec-2001changed wind gen pot againstn2oncley
56 Thu 06-Dec-2001checked Rsw.outtoweroncley
55 Thu 06-Dec-2001guy wire tensiontoweroncley
54 Wed 05-Dec-2001TC data from 21Xtoweroncley
53 Wed 05-Dec-2001precip gauge okaytoweroncley
52 Wed 05-Dec-2001soil Qstn2oncley
51 Wed 05-Dec-2001site visittoweroncley
50 Wed 05-Dec-2001co2 calibrationtoweroncley
49 Tue 04-Dec-2001rainrtoweroncley
48 Tue 04-Dec-2001new snowoncley
47 Tue 04-Dec-2001second mouse trappedbaseoncley
46 Tue 04-Dec-2001added stool!toweroncley
45 Tue 04-Dec-20015m thermocouple fixedtoweroncley
44 Tue 04-Dec-2001Pressure errorall stnsoncley1
43 Tue 04-Dec-2001soil moistureall stnsoncley
42 Tue 04-Dec-2001marigold downtoweroncley
41 Tue 04-Dec-2001Sean's thermocouple commentstoweroncley
40 Tue 04-Dec-2001depth gauge checktoweroncley
39 Tue 04-Dec-2001rain gauge caltoweroncley
38 Tue 04-Dec-2001cal gas valuestoweroncley
37 Mon 03-Dec-2001covars restartedbaseoncley
36 Mon 03-Dec-2001raised breezecom antennatoweroncley
35 Mon 03-Dec-2001wind generator pot turnedstn2oncley
34 Mon 03-Dec-2001unknown crash of marigold and cosmostoweroncley
33 Mon 03-Dec-2001cal gas amountstoweroncley
32 Mon 03-Dec-2001task listall stnsoncley
31 Mon 03-Dec-2001miscstn2oncley
30 Mon 03-Dec-2001last value (20m)toweroncley
29 Mon 03-Dec-2001daily statusall stnsoncley
28 Sun 02-Dec-2001bare sonic moved downstn2oncley
27 Sun 02-Dec-2001Controlling UPSbasemaclean
26 Sun 02-Dec-2001data network configurationall stnsmaclean
25 Sun 02-Dec-2001Dec 2 morning statusall stnsmaclean
24 Sat 01-Dec-2001increased sampling rate on 20m sonictowermaclean
23 Sat 01-Dec-2001licor CO2 cal gas sequencetowermaclean
22 Sat 01-Dec-2001Operational!all stnsmaclean
21 Sat 01-Dec-2001Bare site sonic boom angletowermaclean
20 Sat 01-Dec-2001Prop boom anglestowermaclean
19 Sat 01-Dec-2001Tower sonic boom anglestowermaclean
18 Sat 01-Dec-2001Swapped ironics card in marigoldtowermaclean
16 Fri 30-Nov-2001bare stuffstn2militzer
15 Sat 05-Aug-2000Soil water gravimetric dataall stnsdelany
14 Thu 29-Nov-2001statusall stnsoncley
13 Thu 29-Nov-2001hourly.csh, daily.csh, and SSHbasemaclean
12 Thu 29-Nov-2001ISP and LAN-modembasemaclean
11 Thu 29-Nov-2001Ironics failures & light bulb heaterstowermaclean
10 Thu 29-Nov-2001Tower heightstoweroncley
9 Wed 28-Nov-2001statustoweroncley
8 Tue 20-Nov-2001manual control noteall stnssemmer
7 Thu 15-Nov-2001manual control of TTL portstowermaclean
6 Thu 15-Nov-2001UPS softwarebasemaclean
5 Sat 10-Nov-2001Breezecom settingsall stnsmaclean
4 Sat 10-Nov-2001Adam/network testingall stnsmaclean
3 Sat 10-Nov-2001Adam VME serial numbersbasemaclean
2 Fri 09-Nov-2001Archive size and schemebasemaclean
1 Fri 09-Nov-2001Logbook creatednonemaclean


180: Adams, Site tower, Wed 03-Apr-2002 09:52:49 MST, Teardown
The trusty adams were turned off at 15:10 MST, April 1.

179: Soil, Site tower, Wed 03-Apr-2002 09:50:22 MST, Tower soil samples
The soil samples taken on March 27 were weighed.  Results are in entry #15.
178: Radiometers, Site all stns, Wed 27-Mar-2002 09:41:17 MST, Radiometers and soil sensors removed from all sites
A special trip was made to FLOSS to recover the radiometers and soil probes
Tony, Charlie and Pierce removed the sensors at Bare, Sage and then Tower

(Also the EVE boxes were removed at the bare and sage sites - so no
data from those sites after 15:00 MST March 26.   GDM)

Bare 12:15
Ground not frozen and soil sensors easily uprooted
Net Q97035
Fluxplate H993563
Tsoil NCAR 0001
Soil moisture 013
Soil sensors placed in Netrad box

Campbell 13527

Sage 12:30 
The up looking Pyrg and K&Z were each shadowed using Steve O's wand, for
three successive periods of 30 s at ~ 12:40
Net Q94125
Flux
SoilT
SoilQ

Campbell 11173

Tower 13:30
The up looking Pyrg and K&Z were each shadowed using Steve O's wand, for
three succesive 30 s periods, for three 10 s periods and for one 60s period


Net Q99258
Flux
SoilT
SoilQ

Campbell 11174

The CO2 cal gases: 
#6 Local 0/0
#5 Hical 700/20


177: Soil, Site tower, Wed 27-Mar-2002 09:36:40 MST, Soil core
Aliquots were in oven. These were matched with the soil samples for logbook#171
The aliquots were weighed and the results entered in Lobgook # 15

A soil sample was taken using a spade, placed in a plastic bag and at the base was separated into 0-2, 2-5 and 5-10 cm triple aliquots. These were weighed freshand put in oven for weighing on teardown.
The data is in # 15
176: Kryptons, Site , Mon 18-Mar-2002 08:38:58 MST, Kryptons bad at PAM sutes
The kryptons appear to be suffering from
low temperature problems at the PAM sites.
This has been noted before. The actual cause
has not been determined (As far as I know!)

175: PropVanes, Site tower, Sun 17-Mar-2002 11:32:59 MST, 20m prop down again
The 20 meter prop failed at 13:30 MST March 16.
174: Status, Site , Thu 14-Mar-2002 12:39:36 MST, Daily status, 3/14/02
FLOSS daily status, 03/14/02

This status is based on QC web plots.

Looked at past 48 hours of data
Webcam shows snow was melting till last night. Now
there is a fresh cover of snow.

Meteorology
-----------
Prop Spd and Dir:	ok, prop at 20m replaced yesterday
T/RH (ISFF):		ok, 5 meter T questionable from 3/13/02 17:00 to 5:00
Pressure:		ok
Rainr:			ok, activity throughout last night
Depth:			ok, corresponds with rain activity
Rsw:			ok
Rlw:			ok
Rnet:			ok

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok













173: Kryptons, Site all stns, Wed 13-Mar-2002 16:51:47 MST, rad & kryptons cleaned
Sites 2 first then 3.

172: Radiometers, Site tower, Wed 13-Mar-2002 16:10:37 MST, rad cleaned
Shortly after the krypton cleaning about 30min ago, we also cleaned the pygs and psps (but I forgot net!).

171: Soil, Site tower, Wed 13-Mar-2002 15:52:31 MST, soil samples
Soil Dry weights:

(GDM, April 2: I believe these are actually wet weights of samples taken
March 13, at tower rad stand):
0-2cm	A1:  9.2 B1:  9.6

3-5cm   A2: 24.6 B2: 31.6

5-8cm   A3: 40.2 C2: 39.2

Note added on 27 Mar by TD
On 26 Mar there were aliquots in the ovenI beleive that these correspond
to a core taken at tower on 1 Mar

Note added on April 2 by GDM:  I believe the above note is wrong.
Those weights correspond to a sheet found in the trailer,
in Sandy's handwriting, for samples taken on March 13, 3:30 pm, which
is documented in #168, 13 March, as a sample taken near tower
rad stand.



170: CO2, Site tower, Wed 13-Mar-2002 15:51:50 MST, cal gases
Cal gas amounts:
#5: 1050
other: 710

169: Kryptons, Site tower, Wed 13-Mar-2002 15:51:25 MST, krytpons cleaned
Just cleaned kryptons at 10m and 2m on the tower.

168: PropVanes, Site tower, Wed 13-Mar-2002 14:00:40 MST, several things
Swapped out bad prop at 20m with "spare".  Used the same prop (pitch=306) and reset the boom angle to 177.5 deg as the prop we pulled.

At the same time, cleaned the krypton at 20m.

Also took a core sample near rad stand.

167: Status, Site , Tue 12-Mar-2002 08:28:27 MST, Daily status, 3/12/02
FLOSS daily status, 03/12/02

This status is based on QC web plots.

Looked at past 48 hours of data
Webcam shows snow is melting

Meteorology
-----------
Prop Spd and Dir:	prop at 20m out
T/RH (ISFF):		ok
Pressure:		ok
Rainr:			ok
Depth:			ok
Rsw:			ok
Rlw:			ok
Rnet:			ok

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok












166: Adams, Site tower, Sat 09-Mar-2002 11:01:21 MST, cosmos analog channels flatlined
All analog channels on cosmos made a step change to a constant
value of 1.845 volts at 02:01:30.8 MST, March 9, and stayed there.

1.845 V was the value of channel 102 (co2mr) just before
the change - so it looks like somehow the AD or the controlling
software got fixated on that value.

The co2 cal gas valves switching occured at 02:00, 02:02 and 02:04.
Because they were not coincident, I don't think this problem is 
related to the control of the gas valves.  

rlogging into cosmos showed nothing wrong - though you can't find
out much about the datel A/D from the matrix card.  There isn't
a datel process - just an interrupt function.

An mxreset of cosmos at about 10:45 MST has apparently fixed the problem.

I can't recall seeing this before.




165: PropVanes, Site , Fri 08-Mar-2002 13:24:24 MST, 20 meter prop out
The 20 meter prop is questionable.
This occurred at ~12:00, noon, local.


164: Radiometers, Site stn3, Mon 04-Mar-2002 11:29:04 MST, 4-component aspiration fans disconnected 3/2
4-Component FANS Disconnected:  3/2/02 ~16:30z
to reduce the power consumption.

The dome ventillator fans were pulling too much current
when the temperature was dropping too low.
There will probably be no measurable effect in doing
this because of minimal dust and ample winds in
North Park during this time of the year.

Note this was also entered in the 'station outage'
log entry.
163: Status, Site stn3, Mon 04-Mar-2002 10:10:12 MST, Sta 3 Outages 3/2, 3/3
Station 3 Power Outages

It appears Sage site is getting cold feet.....

---------------------------------

Sta3 - Sage Outage:
3/2/02 ~11:30 - 15:00z
Low voltage disconnect did it's thing.
Station recovered by itself with adequate solar charging,
no wind charging at the moment.

4-Component FANS Disconnected:  ~16:30z
to reduce the load.  They're still jumping up and down in
power consumption when the temp. drops too low.

---------------------------------

Sta3 - Sage Outage:
3/3/02 ~9:15 - 14:45z
Low voltage disconnect did it's thing again.
Station recovered by itself with adequate solar charging.
At about local noon, the winds picked up and the wind
generator was providing charging through the night.



162: Soil, Site tower, Fri 01-Mar-2002 09:41:06 MST, Soil core at tower
Took soil core at tower and separated into 0-2,2-5 and 5-10.
See logbook #15? for data (also my paper logbook)
161: Kryptons, Site stn3, Fri 01-Mar-2002 09:37:19 MST, Replace Kr at Sage
Inspected Sage site. All vent + asp fans OK
Kr #1133 was oriented with the receiver pointing to the S.
Kr 1133 was removed.
Kr 1395 was installed with receiver pointing to the N (wire S of attachment)
Had trouble sucuring the device but got screwws started and then used panduits.

160: Status, Site tower, Wed 27-Feb-2002 16:16:33 MST, Guy tension/chem shelter
Checked guy tensions: see record # 109
Chem shelter moved at 14:30 to position near the entry gate, to remove it
from the upwind fetch of the precip guage.
159: PropVanes, Site tower, Wed 27-Feb-2002 14:32:19 MST, 20 m propvane reinstalled
At 14:15 the 20 m propvane was reinstalled.
Propvane 005 with prop FLOSS 5, #56159

Back at trailer the wind dir/spd dispay showed reasonable values.
158: Snow_Precip, Site tower, Wed 27-Feb-2002 14:29:40 MST, Rain guage inspection
At 13:45 visited tower and inspected precipitation guage.
The top of the cylinder is level, E-W, N-S.
The wind baffles are installed with the concave side outermost.
The stakes are firm
157: Soil, Site , Wed 27-Feb-2002 10:54:05 MST, Soil cores at Sage and Bare
At 9:15 visted Sage. Asp and vent fans OK. Brushed away ~8cm powder snow and 
took a soil core. Back in trailer subdivided into 0-2cm, 2-5cm, and 5-10cm.

At 9:15 visited Bare. Asp fan OK. Brushed away  ~5cm powder snow and took a
csoil core. Back in trailer subdivided into 0-2cm, 2-5cm and 5-10cm.

All the soil data reported in # 15

156: CO2, Site tower, Tue 26-Feb-2002 14:31:04 MST, Removed Licor 7500
Between 14:00 and 14:15 we removed the Licor 7500, brought it back to the 
trailer and packed it into its box
155: PropVanes, Site tower, Mon 25-Feb-2002 15:07:49 MST, Reinstalled 1 m propvane
Atempted to reinstall propvane # 004 /prop 56153 but spinning prop broke. Blood! 
At 14:45 installed propvane #007 / prop 53560

Now there is only one spare propvane with matching prop
154: Camera, Site tower, Fri 22-Feb-2002 13:34:32 MST, Axis webcam settings
Here are the current settings in the Axis 2420 camera:

MAC address: 00:40:8C:56:E0:E0
Image - General
  Appearance
    Images: single
    Resolution: 352x240 ~13Kb
  Tuning:
    Compression: Low
    Brightness: 10  (0..15)
    White Balance: Outdoor
    Color level: 6  (0..15)

  Heading:
    Text:  FLOSS
    Date: yes
    Time: 24h
  Default viewer:
    Java Applet

Image - Advanced
  Appearance:
    Sharpness: 166  0..255
  DC Iris:
    Exposure Control: DC-Iris
    Shutter Speed: 1/60
    Gain: 22dB   (10,16,22,28)
    ALC Level: 0

Image - Focus
  Assistant disabled

Network - TCP/IP
  Enable DHCP

  Internet addr: 192.168.10.40  (done via DHCP)
  Subnet mask: 255.255.255.0    (ditto)
  Default router: 192.168.10.1  (ditto

  HostName: webcam

  Domain Name: atd.ucar.edu
  Primary DNS: blank
  Secondary: blank


  Select Media: auto-negotiate
  Max Bandwidth: 1.0 MBit/s
  HTTP Port Number: 80


Network - SMTP
  Primary Mail server: blank
  Secondary Mail server: blank

Network Nofification:
  Http: disabled

  Smtp: disabled

  Ftp: disabled

COMM ports:
  RS232:
    Modem
    Baud rate: 9600
  RS485:
    none


Modem General:
  defaults

System - Date and Time

  New Camera time:
    Synchronize with NTP
      IP: 192.168.10.1
      TimeZone: GMT-07 Arizona, Mountain Time

System - Users
  root: ADV
  aster: ADV


Applications
  Operations - Scheduler
    Sequential Mode
    Primary Time Enabled: yes
      Restricted between
        Start: 05 00
        Stop:  20 00
        Mon-Sun
      Primary Image Freq:
        Every 15 mintues
        Regardless of input

  Sequential Operation - Upload
    Upload via: FTP, using: Network
    Remote Host: 192.168.10.1
    User name: aster
    password:

    Use Passive Mode: no
    FTP Port number: 21

    Image file:
      size: 352x240
      upload path: /backup/aster/projects/FLOSS01/webcam
      base file name: floss

      Overwrite: no
      Date/time suffix: yes
      Sequence # up to def max: no
      Sequence # up to spec max: no blank
  Application: enabled


Motion detection:
  not used

Layout - General:

  Enable personal settings: no

 

153: Kryptons, Site all stns, Fri 22-Feb-2002 11:25:17 MST, Kr and Rad clean
At 10:50 visited Sage and cleaned Kr and all radiometers
At 11:00 visited Bare and cleaned Kr and netrad
At 11:10 visited tower and cleaned Kr.2m and all radiometers

Note in no case was there any indication of dirt on the domes.
152: PropVanes, Site tower, Fri 22-Feb-2002 11:20:47 MST, Remove 1m propvane
At 09:35 I removed the 1m prop vane to take to NCAR for repairs
151: Status, Site tower, Thu 21-Feb-2002 14:54:57 MST, Guy wire tension adjustment
Between, 13:30 and 14:30 I adjusted the tower guy tensions.
First I monitored the 12 individual guys. Then I increased the tension on the 
lowest value guy by one notch. This increased the value for that guy by a 
several hundred pounds. It also modulated the tensions on the other guys.
I monitored all the guys again and then increaded the tension on the new
lowest value guy. I reiterated three times until all tensions were above 400lbs.

See logbook entry 109 for the record.

150: Snow_Precip, Site tower, Thu 21-Feb-2002 12:51:37 MST, Ultrasonic snow depth check
Place plastic toboggan atop small step ladder positioned under the ultrasonic
detector from 10:11 to 10:27. Surface 86 cm below sensor

Placed chem shelter door atop toppled step ladder from 10:28 to 10:39. Surface
1.7 m below sensor.
149: CO2, Site tower, Thu 21-Feb-2002 12:48:36 MST, Cal gas pressure log for #5 & #6
At 10:15 Thu, 21 Feb, checked cal gas pressures
Lo-cal  cyl #6 1700/20
Hi-cal  cyl #5 1600/20
 
compared to 1750/20
            1690/20
 when installed yesterday, Wed, 20 Feb afternoon

On Fri at 09:30
Lo-cal Cyl # 6 1670/20
Hi-cal cyl # 5 1600/20

On Mon,25 Feb, at 14:20
Lo-cal cyl # 6 1480/20
Hi-cal cyl # 5 1480/20

On Tue, 26 Feb at 14:00
Lo-cal cyl # 6 1400/20
Hi-cal cyl # 5 1400/20

On Wed, 27 Feb at 14:30
Lo-cal cyl # 6 1400/20
Hi-cal cyl # 5 1400/20                 check whether solenoids are operating
148: Radiometers, Site all stns, Thu 21-Feb-2002 12:38:40 MST, Radiometer visits
At 09:45 visited Sage. No frost on radiometers but some on frost detector and 
on top dome of net radiometer. Cleaned frost from frost detector and net 
radiometer. Handkerchief placed over frost detector for 10 minutes.
All fans operating. Power dissapation resister cold.

At 10:08 visited tower site. No frost on radiometers nor frost detector. Net
radiometer had frost on upper dome. Covered uplooking frost detector from 10:09
until 10:30. Covered downlooking frost detector from 10:30 til 10:40.

147: Soil, Site tower, Wed 20-Feb-2002 16:38:18 MST, Soil moisture profile
When taking the soil cores earlier in the day it was noted that the moisture
was mainly in the top few cm.
At 15:15 a 16 cm deep pit was made at the tower soil site. It was dug with a 
shovel and squared off with a trowel. Samples of 0-2cm, 2-4cm, 4-8cm and 8-16cm
were extracted and loaded into tupperware containers. These were processed in 
the trailer. The 0-2cm sample had substantial amount of grass, the next two 
sample had a fair amount of root fiber, and the deepest sample was mostly dry 
dirt.

             Fresh      Dry
0-2cm  DI   5.8g       3.6g
       DII  4.7g       2.8g
       DIII 5.9g       3.6g

             Fresh      Dry
2-4cm  EI   14.9g      11.9g
       EII  17.3g      13.7g
       EIII 19.0g      15.2g

             Fresh      Dry
4-8cm  FI   19.2g      17.1g
       FII  18.7g      16.7g
       FIII 20.4g      18.3g

             Fresh      Dry
8-16cm GI   26.3g      24.9g
       GII  23.0g      21.8g
       GIII 29.0g      27.4g

The soil samples were taken and weighed on Wed, 20 Feb''02.
They were dried in the oven at 110C and weighed on Thu and again on Fri.
This was to check that drying was complete after a single day.
146: Soil, Site all stns, Wed 20-Feb-2002 16:30:14 MST, Soil samples at sage, bare and tower
At 10:00 0-5cm soil core taken at sage. Several cm of fresh snow first removed.
      Fresh    Dry
AI   19.4g    16.5g
AII  20.4g    17.4g
AIII 23.4g    19.7g

At 10:30 0-5cm soil core taken at bare. Several cm of fresh snow first removed.
      Fresh    Dry
BI   19.6g    16.7g
BII  23.9g    20.1g
BIII 27.1g    22.7g

At 11:00 0-5cm soil core taken at tower A cm of fresh snow first removed.
      Fresh    Dry
CI   17.6g    13.2g
CII  17.0g    12.7g
CIII 16.6g    12.0g

The soil samples were taken  on Wed, 20 Feb''02, and weighed
They were cooked in the oven at 110C and reweighed on Thu. To check the 
completeness of drying they were left in the oven until Fri AM and reweighed.
The Thu and Fri weights were identical.
145: CO2, Site tower, Wed 20-Feb-2002 16:22:08 MST, Change cylinders, check flow rates
At 14:00 checked intake flow of CO2 flux system = 4 lpm. 
This low value prbably because of the progressive loading of the intake filter.

Changed cylinders to
 Hi-cal cylinder # 5 Start at 1690/20
 Lo-cal cylinder # 6 Start at 1750/20

Adjusted bypass valve to make intake flow = 7.5 lpm

Cmmunication with PAM base, ttl_on/off cosmos 0/1 then ttl_return cosmos 0/1
Adjusted lo-cal flow valve to yield an excess lo-cal flow of 1 lpm
Adjusted hi-cal flow valve to yield an excess hi-cal flow of 1 lpm




144: PropVanes, Site tower, Tue 19-Feb-2002 18:06:15 MST,
An attempt was made to fix the 0.5m propvane by swapping parts: 
The end result was to have a valid speed but no direction.
The components are #4 chopper and prop and vane and electronics and
#7 encoder.
Since the encoder is different, the boom azimuth will have to be reset (but
direction isn't working now anyway).

Summary:
Either chopper read 0 through #7's electronics.
Either encoder read constant through #4's electronics.
Thus, the problem has to be with the electronics? 

The 20m propvane was brought down to be taken to NCAR.

143: Kryptons, Site stn2, Tue 19-Feb-2002 14:27:43 MST, new krypton
at bare, just swapped krypton 1101 for 1392, again to attempt to solve the
cold temperature problem.

142: Kryptons, Site stn3, Tue 19-Feb-2002 14:25:28 MST, new krypton
at sage, we just swapped 1133 for 1395 in hopes of solving the cold temperature
dropout problem

I also checked the fence battery which is now 14.5V connected through the 
solar panel, so I think all is well.

I also note the serial numbers:
csat 538
trh 703

141: Soil, Site tower, Mon 18-Feb-2002 17:09:10 MST, soil samples
Soil samples taken at main tower about 16:30 today.  Someone hadn't put the 
brass rings back in the case, so I didn't have them.  Thus, I've processed 
this sample as one chunk, approx 3.75" thick.  I made 3 alloquots from it:
Wet:		Dry:
A1 36.3		33.3
A2 26.7		24.3
A3 28.8		25.4
Also, the total weight of the sample was 265.6, so a crude bulk density could
be calculated.
(We need a small metric ruler here!)


140: Thermocouples, Site tower, Mon 18-Feb-2002 15:48:29 MST, thermocouples installed
Thermocouples are now in all positions.  These are the "thick wire" versions
from Sean.  This done ~13:00.

139: Kryptons, Site all stns, Mon 18-Feb-2002 15:47:14 MST, kryptons cleaned
All kryptons have now been cleaned, starting with the top of the main tower at
about 13:00 and progressing down, the sage ~1500 and bare ~1530.

138: Radiometers, Site all stns, Mon 18-Feb-2002 15:46:12 MST, rads cleaned
All rads have now been cleaned.
13:55 local - main tower
~15:00 local sage site
~15:15 local bare site

137: Power, Site all stns, Mon 18-Feb-2002 15:43:45 MST, power system inventory
PAM charge controllers (inventory for IHOP):

bare:
main box: Flexcharge NC25A-12 and Sunsaver 20 in cooler + 5 batteries
fence charger: Sunsaver 20 in old-style box + 1 battery


sage:
main box: Flexcharge NC25A-12 and Sunsaver 20 in cooler + 4 batteries
fence charger: Prostar 30 in old-style box + 1 battery

136: CO2, Site tower, Mon 18-Feb-2002 15:41:11 MST, low? CO2 cylinder changed
I just swapped the nearest cylinder on the closed-path CO2 system, since it was
out.  We're now using cylinder #5, which had about 1900 lbs of pressure 
initially.  The other cylinder still has about 500 lbs in it, so I didn't swap
it.

P.S. The cylinder status now is:
cyl #2 - now empty
cyl #6 - not yet used -- ready to swap with #4 when it dies.
cyl #4 - still had 410 lbs as of 4pm yesterday, connected to system
cyl #5 - lost 200 lbs yesterday, so only 1700 lbs as of 4 pm yesterday.
	it is now shut off, but connected to the system.


135: PropVanes, Site tower, Sun 17-Feb-2002 07:49:58 MST, 20m prop speed decoder out
The 20 meter prop speed is zero starting at Feb 14 14:00 MST

The prop at 20 meters has been replaced before, see entry #103

134: Camera, Site tower, Sun 17-Feb-2002 07:42:44 MST, Lost webcam pictures Feb14-16
Due to human error (who me?) we've lost all web cam photos for Feb 14-16.
133: Status, Site all stns, Tue 12-Feb-2002 14:44:22 MST,
FLOSS status, 2002 Feb 11

NOTE: I have just started the data minding task


Webcam:


        Meteorology
        -----------
        Prop Spd/ Dir:  ok     
        T/RH (ISFF):    ok
        Pressure:       ok
        Rainr:



        Depth:          ok value varies
        Rsw:            ok the albedo at the tower is ~60%, at sage only 30%
        Rlw:            ok
        Rnet:           ok different albedos

        Soil
        ----

        Tsoil:          ok, tower still noticably warmer
        Gsoil:          ok, sage still has significantly less flux
        Qsoil:          ok, sage still driest


                        Sonics
        ------
        flags/diag              ok, flag and diag hits on tower 2m.  This appears
                        correlated with temps below -20degC.
        spd, dir, w:            ok the difference plot now fixed        
        variances:              ok the differnce plot now fixed
        u*, sigma_w/u*, w'tc':  ok


        Scalars
        -------

        H2O             ok, The open path Licor is strange beast, probably a calib.
                        issue.  It apparently has inverse trend and
                        gain factor of about 12x compared with kryptons.
                        Sage/Bare still have irregularly high signals
                        at times and these correlate with very low temps
                        going below ~-22degC. 
        co2:            ok still showing calibration peaks, ie still got cal gas
        w'h2o':         ok
        w'co2':         ok


        Status
        ------
        PAM power
        X/Ylevel
        Ice    

        <><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><>



132: Status, Site , Sat 09-Feb-2002 10:43:20 MST, Weekly Status, 2-9 Feb 2002
FLOSS status, 2/9/02

NOTE: I have been looking at data for the last few days.
      This is a combined simplistic status report for the
      last week as opposed to a daily report.

Webcam still has low-resolution, grainy images but appears
	to show good snow cover although it's impossible to
	tell if it's changed much over the period.

Meteorology
-----------
Prop Spd/ Dir:	ok
T/RH (ISFF):	ok
Pressure:	ok, offset in tower analog transducer, calib.?
Rainr:		ok?, many tips with diurnal cycle apparently occuring
			below -10degC.  The sum for 7 days would have been
			1.58" if we can believe that.  There was some snow 'in
			the high country' reported this week.
Depth:		ok, still has diurnal cycle, but has gone up about 10mm
Rsw:		ok, sage.out not as high as tower.out
Rlw:		ok
Rnet:		ok, tower and bare very low values, maybe snow on top
			(but hasn't melted...?), or dirty, etc.  Sage ok.

Soil
----
Tsoil:		ok, tower still noticably warmer (1-deg), cal/instl.?
Gsoil:		ok, sage still has significantly less flux
Qsoil:		ok, sage still driest

Sonics
------
flags/diag		ok, flag and diag hits on tower 2m.  This appears
			correlated with temps below -20degC.
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
h2o:		ok, The open path Licor is strange beast, probably a calib.
			issue.  It apparently has inverse trend and
			gain factor of about 12x compared with kryptons.
		Sage/Bare with serializer still have irregularly high signals
			at times and these correlate with very low temps
			going below ~-22degC.  Could be a calibration issue
			on the serializer a/d.   Only raw data counts reported
			so we're not taking advantage of integrated sensor
			approach.
co2:		ok
w'h2o':		ok
w'co2':		ok


Status
------
PAM power	ok, both stations batt staying up nicely
X/Ylevel	ok, diurnal cycle on tower sensor, X as high as -1deg
Ice		ok?  Got a hit on 5th at about 20:00-20:30 so probably real.














131: Camera, Site tower, Wed 06-Feb-2002 15:52:27 MST, network camera
I figured out how to talk to the axis network camera from Boulder.

I made these changes (the gain change was recommended by tech support
at axis.com):

  Raised gain from 10dB to 22dB  (other choices are 10,16,22,48)
  Changed brightness from 8 to 10  (choices are 0...15)

  Changed resolution of FTP'd photos from 704x480 to 352x240 in
  order to avoid the file truncation problem.




130: Status, Site , Fri 25-Jan-2002 07:18:41 MST, Daily status, 1/25/02
FLOSS daily status, 1/25/02

Looked at past 48 hours of data
Webcam shows snowfall over surface

Meteorology
-----------
Prop Spd and Dir:	ok, winds around 200 degrees
T/RH (ISFF):		ok
Pressure:		ok
Rainr:			ok, slow down in tips
Depth:			ok, has gone up, maybe snow melt
Rsw:			ok, sage.out not as high as tower.out
Rlw:			ok
Rnet:			ok, tower and bare low values, maybe snow on top?

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok











129: Status, Site , Thu 24-Jan-2002 07:18:48 MST, Daily status, 1/24/02
FLOSS daily status, 1/24/02

Looked at past 48 hours of data
Webcam shows snowfall over surface

Meteorology
-----------
Prop Spd and Dir:	ok, winds around 200 degrees
T/RH (ISFF):		ok
Pressure:		ok
Rainr:			ok, a lot of tips during day 23
Depth:			ok, continued to move down
Rsw:			ok, sage.out not as high as tower.out
Rlw:			ok
Rnet:			ok

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok










128: Status, Site , Wed 23-Jan-2002 08:22:35 MST, Daily status, 1/23/02
FLOSS daily status, 1/23/02

Looked at past 48 hours of data
Webcam shows showfall started late yesterday

Meteorology
-----------
Prop Spd and Dir:	ok, winds 150-200 degrees
T/RH (ISFF):		ok
Pressure:		ok
Rainr:			ok, a lot of tips from last night to this morning
Depth:			ok, change started at midnight. Corresponds with
			    rainr info. Webcam confirms snowfall
Rsw:			ok, sage.out not as high as tower.out
Rlw:			ok
Rnet:			ok

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok









127: Status, Site , Wed 23-Jan-2002 08:19:15 MST, Daily status, 1/22/02
FLOSS daily status, 1/22/02

Looked at past 72 hours of data


Meteorology
-----------
Prop Spd and Dir:	ok, winds 150-250 degrees
T/RH (ISFF):		ok
Pressure:		ok
Rainr:			ok, some tips
Depth:			ok, spiking
Rsw:			ok
Rlw:			ok
Rnet:			ok

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok






126: Status, Site , Tue 22-Jan-2002 10:25:01 MST, Daily status, 1/18/02
FLOSS daily status, 1/18/02

Looked at past 72 hours of data
Comments:  One has to wonder about the continuous sporadic occurrence of
           single tips on the rain gauge.
           Entered into logbook on 1/22/02, since logbook not available 
           on 1/18/02.

Meteorology
-----------
Prop Spd and Dir:	ok, winds mostly 180-270 degrees
T/RH (ISFF):		ok; did not examine thermocouple data
Pressure:		tower 0.5-3 mb higher than bare/sage;
			  sage 0.1 mb higher than bare
Rainr:			ok(?); netcdf file has units of tips/60
Depth:			"snow depth" ~ 2050 mm; difficult to interpret
Rsw:			ok
Rlw:			Rlw.in.tower > Rlw.in.sage by ~5 W/m^2
Rnet:			ok; Rnet < Rsum by 5-10%

Soil
----
Tsoil:			sage has a little larger diurnal cycle than
			  tower or bare
Gsoil:			sage has smallest diurnal cycle, bare greatest
Qsoil:			sage driest, bare wettest; tower very constant

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			not much correlation evident between 
			  w'co2'(LI7500) and w'co2mr'(LI6251);
  			  LI7500 has diurnal cycle; LI6251 much noisier
co2:			good correlation between co2(LI7500) and 
                          co2mr(LI6251); both are anti-correlated with
			  temperature


125: Status, Site , Tue 22-Jan-2002 10:22:56 MST, Daily status, 1/17/02
FLOSS daily status, 1/17/02

Looked at past 72 hours of data
Comments:  Fixed "overrange" problem for KH2O @ sage; we were using the
wrong calibration; Gordon reran covars.
Entered into logbook on 1/22/02, since logbook was not available on
1/17/02.

Meteorology
-----------
Prop Spd and Dir:	ok, winds mostly 180-270 degrees
T/RH (ISFF):		ok; did not examine thermocouple data
Pressure:		tower 0.5-3 mb higher than bare/sage;
			  sage 0.1 mb higher than bare
Rainr:			ok; netcdf file has units of tips/60
Depth:			"snow depth" ~ 2050 mm; difficult to interpret
Rsw:			ok
Rlw:			Rlw.in.tower > Rlw.in.sage by ~5 W/m^2
Rnet:			ok; Rnet < Rsum by 5-10%

Soil
----
Tsoil:			sage has a little larger diurnal cycle than
			  tower or bare
Gsoil:			sage has smallest diurnal cycle, bare greatest
Qsoil:			sage driest, bare wettest; tower very constant

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			not much correlation evident between 
			  w'co2'(LI7500) and w'co2mr'(LI6251);
  			  LI7500 has diurnal cycle; LI6251 much noisier
co2:			good correlation between co2(LI7500) and 
                          co2mr(LI6251); both are anti-correlated with
			  temperature


124: , Site , Fri 18-Jan-2002 10:06:45 MST, RAD domes cleaned at tower and sage
NOt much visable contaminents on domes.

Note: sage was cleaned late yesterday.



123: , Site , Fri 18-Jan-2002 10:04:53 MST, CAL GAS readings
CAL GAS readings:
cal bottle 2 - 1150
cal bottle ? - 1400

I could not find the identification for the one gas bottle

122: , Site , Fri 18-Jan-2002 10:03:55 MST, trime at tower
Trime readings at tower. 3 readings taken 1 minute apart.
1 - 10.1
2 - 10.3
3 - 10.1

121: Status, Site tower, Fri 18-Jan-2002 08:40:35 MST, Installed web cam
We installed the axis web camera on the tower just above the 20m prop vane.


Right now this is the configuration for storing photos:

Between 05:00 and 20:00 MST it will store 704x640 images
on /backup/aster/projects/FLOSS01/webcam, every 15 minutes.


There is a problem however: it is truncating the images, only
56000 bytes are being ftp'd, and so we're only getting about
2/3rds of each image.  I've sent an email inquiry to axis.com.







120: Kryptons, Site stn2, Fri 18-Jan-2002 08:28:56 MST, Added h2oV.bare
Last night we changed the setup at stn2 (bare site) to sample the
krypton voltage through the Eve A2D, in addition to the serializer.

The new variable is h2oV.bare - it is being covar'd into the netcdf files.


Last night, at least, the voltage reported by the serializer agrees
with h2oV - they were both around 2 V.


Note that the serializer voltage as reported with the Eve "KRYPTON"
command on the console appears to be wrong.  

The q_kh2o_serializer cal function converts serializer counts to volts with

  volts = counts * 5 / 32767

On a big-endian system (eve, and suns) you must swap the bytes
of the counts before the conversion.  Eve is either not swapping
bytes or is using some other counts->volts conversion.

**** Note this this problem only effects the display on the eve
console of krypton voltage from the serializer - it does not
effect the values computed by the base system. ****


119: Soil, Site stn3, Thu 17-Jan-2002 17:46:48 MST, Trime at sage
Two readings were taken approximately 5 minutes apart.
1 - 7.6  
2 - 8.2


118: Soil, Site all stns, Thu 17-Jan-2002 17:44:18 MST, soil samples
soil samples:  

wet
	tower		sage
	a1-14.3		b1-29.7
	a2-16.3		b2-31.0
	a3-16.3		b3-31.3
dry (weighed 2/18/02 by spo)
        a1-15.9		b1-25.7
	a2-17.4		b2-26.4
	a3-18.0		b3-27.0
(how can the an samples be higher???)


117: CO2, Site tower, Thu 17-Jan-2002 13:17:37 MST, added tubing to cal gas lines

116: Soil, Site all stns, Thu 17-Jan-2002 10:56:41 MST, soil sample from Tony's visit
These are the dry numbers. Could not find enter for
the wet numbers

A1 - 5.8
A2 - 7.1  
A3 - 9.0

B1 - 16.0
B2 - 15.0
B3 - 15.8
  
115: Kryptons, Site stn3, Wed 16-Jan-2002 19:36:40 MST, Wrong cal used for krypton at sage site
Steve Oncley found the problem with the krypton at the sage site:

  prep.config is using cal_files/krypton1392 for both s2 and s3.
  s3 should have krypton1395.  s3 has been going below 0 since 
  its voltage has been about 3.8 V and Vo for 1392 was 3.6V. 
  Vo for 1395 is 4.9V.


This has now been corrected, and the netcdf data for the sage h2o
variable are being re-computed from the present back to Dec 17,
when it was installed.




114: Status, Site , Wed 16-Jan-2002 12:19:19 MST, Daily status, 1/16/02
FLOSS daily status, 1/16/02

Looked at past 72 hours of data
Comments:  snow depth difficult to interpret.  KH2O @ sage overrange.

Meteorology
-----------
Prop Spd and Dir:	ok, winds mostly 180-270 degrees
T/RH (ISFF):		ok; did not examine thermocouple data
Pressure:		tower 0.5-3 mb higher than bare/sage;
			  sage 0.1 mb higher than bare
Rainr:			ok; netcdf file has units of tips/60
Depth:			"snow depth" ~ 2050 mm; difficult to interpret
Rsw:			ok
Rlw:			Rlw.in.tower > Rlw.in.sage by ~5 W/m^2
Rnet:			ok; Rnet < Rsum by 5-10%

Soil
----
Tsoil:			sage has a little larger diurnal cycle than
			  tower or bare
Gsoil:			sage has smallest diurnal cycle, bare greatest
Qsoil:			sage driest, bare wettest

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			KH2O at sage appears to be over range
 			  LI7500 ok
w'co2':			not much correlation evident between 
			  w'co2'(LI7500) and w'co2mr'(LI6251);
  			  LI7500 has diurnal cycle; LI6251 much noisier
co2:			good correlation between co2(LI7500) and 
                          co2mr(LI6251); both are anti-correlated with
			  temperature


113: Status, Site , Tue 15-Jan-2002 10:24:42 MST, Daily status, 1/15/02
FLOSS daily status, 1/15/02

Looked at past 72 hours of data
Comments:  snow depth difficult to interpret.  KH2O @ sage died.

Meteorology
-----------
Prop Spd and Dir:	ok, winds mostly 180-270 degrees
T/RH (ISFF):		ok; did not examine thermocouple data
Pressure:		tower 0.5-3 mb higher than bare/sage;
			  sage 0.1 mb higher than bare
Rainr:			ok
Depth:			"snow depth" ~ 2050 mm; difficult to interpret
Rsw:			tower & sage only; albedo decreasing at tower
Rlw:			tower & sage only; ok
Rnet:			ok; Rnet < Rsum by 5-10%

Soil
----
Tsoil:			sage has a little larger diurnal cycle than
			  tower or bare
Gsoil:			sage has smallest diurnal cycle, bare greatest
Qsoil:			sage driest, bare wettest

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			KH2O at sage appears to have died around 1000,
			  Jan 14, over 2 hrs since last precip. 
 			  LI7500 ok
w'co2':			not much correlation evident between w'co2' 
			  from LI7500 and w'co2mr' from LI6251;
  			  LI7500 has diurnal cycle; LI6251 much noisier
co2:			good correlation between co2(LI7500) and 
                          co2mr(LI6251); both are anti-correlated with
			  temperature


112: Status, Site , Mon 14-Jan-2002 11:47:17 MST, Daily Status, 1/14/02
FLOSS daily status, 1/14/02

Looked at past 72 hours of data

Meteorology
-----------
Prop Spd and Dir:	ok, winds mostly 180-270 degrees
T/RH (ISFF):		ok; did not examine thermocouple data
Pressure:		tower 0.5-2.5 mb higher than bare/sage;
			  sage 0.1 mb higher than bare
Rainr:			43 mm precip past 72 hours
Depth:			"snow depth" ~ 2050 mm; 
			  not much correlation apparent with precip
Rsw:			tower & sage only; albedo decreasing at tower
Rlw:			tower & sage only; ok
Rnet:			ok

Soil
----
Tsoil:			sage has a little larger diurnal cycle than
			  tower or bare
Gsoil:			sage has smallest diurnal cycle, bare greatest
Qsoil:			sage driest, bare wettest

Sonics
------
spd & dir:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			KH2O ok; correlation coefficient between 
 			  LI7500 and KH2O is 0.97; linear fit is 
			  w'h2o'(LI7500) = w'h2o'(KH2O)*36.9 + 0.0096
w'co2':			not much correlation evident between w'co2' 
			  from LI7500 and w'co2mr' from LI6251
co2:			correlation coefficient between co2(LI7500) and 
                          co2mr (LI6251) is 0.93; 
			  linear fit is co2mr = 33.3*co2mr - 16.3



111: Logistics, Site all stns, Mon 07-Jan-2002 10:33:12 MST, Visit 01/04/2002 Task CheckList Review
====================================================================
			FLOSS Service Visit Tasks
				4-Jan.-02
====================================================================

TASKS:
Done: - Site First:
	Place stool under Snow Depth Gauge SR50 while there.
	Later on checked and exposed thermocouple to atmosphere.

Done: - Tower Guy Wire Tension check
	(update logbook entry 55)
	Confirm they are consistent before climbing...
	(alert Kurt if change by more than 50 lbs)

Done: - Tower Thermocouples
	Install 6-each Thermocouples
	Retrieve dead Thermocouples on lower levels

Done: - Tower Photographs
	near top of tower
	See sample photographs with camera.

Done: - Tower 20m Prop Change/Fix dead

Done: - Tower Clean Kryptons	 (alcohol / Q-tips)

Done: - Tower	Clean Radiometers (dry wipes)

Done: - Tower Cal-Gas
	Change 2 cylinders  (update logbook entry 33)
	check flow rate at CO2 sensor inlet
NOT: 	(create logbook entry on flow rate)
NOT:	check	turn on/off:	'ttl_on cosmos 1'
				'ttl_on cosmos 0'

Done: - Tower Precip Gauge fluid level check
	(drain if within 6" top of the gauge)

Done: - PAMs	Clean Kryptons	 (alcohol / Q-tips)
Done: - PAMs	Clean Radiometers (dry wipes)
NOT:  - PAMs	Find/Retrieve Prototype EVE Power Supply
	Could not find this power supply...bummer...
NOT:  - PAM	Bare site Vbatt monitor bad
	Take replacement bbox next time

NOT:  - Trime soil moisture reading; each site
	(update logbook entry 63)
        Battery dead; recharge and take next time.

Done: - Gravimetric soil sample at each site
	(logbook entry 52; clone logbook entry 62)

Done: - Base	Swap backup disk in aster
	Reboot / Check aster

Done: - Base	Retrieve Licor Manual

Done: - Base	Update LOGBOOK

Done: *** DON'T FORGET: REMOVE STOOL FROM UNDER SR50 ***





110: Kryptons, Site stn3, Fri 04-Jan-2002 16:43:13 MST, PAM Kr cleaned
Cleaned Kr at the two PAM stations at 16-30

109: Logistics, Site tower, Fri 04-Jan-2002 16:22:09 MST, guy wire tension update
Guy wire tensions Update

Note: I re-arranged the listing to hopefully make it easier
to track the slackening.


UpDate	 North-East	 South-East	 South-West	 North-West
	 L    M    H	 L    M    H	 L    M    H	 L    M    H
120601	520  650  550,	690  620  480,	620  450  670,	620  490  630
120901	510  650  550,	670  590  480,	560  430  660,	600  480  600
121801	500  600  520,	630  550  440,	570  400  590,	600  475  600
010402	460  540  490,	580  520  400,	540  370  570,	575  440  590
011702	460  550  450,  590  500  390,  570  370  600,  590  450  590
021802  370  470  400,  470  430  410,  480  330  520,  470  350  510
022102  390  550  400,  500  450  350,  510  330  540,  470  360  510
Tony Delany adjusted tensions, see logbook # 151, 21 Feb''02
        660  540  470,  500  410  540,  550  580  480,  450  420  620
022702  690  550  450,  520  430  480,  600  570  470,  560  440  640




108: Base, Site base, Fri 04-Jan-2002 16:17:35 MST, Swapped Base Disk Drive
01/04/2002 16:10

Removed disk 'ISFF backup #2'
Installed disk 'ISFF backup #1'
Reboot of aster worked perfectly.
107: Snow_Precip, Site tower, Fri 04-Jan-2002 16:15:25 MST, Snow Depth Test: Thermocouple exposed to air
01/04/2002

Placed short step-ladder under snow depth sensor
13:00 installed ladder
16:10 removed,

Note Thermocouple was inside and the cap was on.
It was pulled outside at 16:10 such that it won't
be seen by the sun (ie ~1cm under the box).


106: Radiometers, Site all stns, Fri 04-Jan-2002 15:57:35 MST, Radiometer Domes cleaned
12:45 Radiometers on Tower, all 6 fans OK, Rnet level OK.

15:30 Radiometers on Sage-brush site, all fans OK, Rnet level OK,

On both sites the domes appeared clean before wiping off.

On both sites there was some snow on top of the up-looking radiometers shield.
105: Kryptons, Site tower, Fri 04-Jan-2002 15:56:25 MST, Kryptons Cleaned
01/04/2002

13:35 2m
13:50 10m
14:00 20m

104: Thermocouples, Site tower, Fri 04-Jan-2002 15:53:41 MST, New Thermocouples Installed.
01/04/2002 ~13:30-15:00

I removed the 4 dead thermocouples and installed 6 new ones at
.5,2.0,5.0,2x8.0 and 20.0m
The 'scripts/profiles' script was edited to remove the 'non-installed'
levels so the xtemp display was appropriate.

103: PropVanes, Site tower, Fri 04-Jan-2002 15:50:49 MST, 20m Prop Replaced
01/04/2002 ~15:00
We replaced the prop at 20m which had the speed out.
At 15:50 local I changed the boom angles in the sensor per
the logbook entry.   Prior to that the direction angle was
out by 180-degrees after we put in the new sensor.

W:0005    696    91  2074   50 00I\n\r
W:0005    703    92  2062   50 00V\n\r
W:0005    703    92  2053   50 00V\n\r
W:0005    703    92  2053   50 00V\n\r
W:0005    703    92  2053   50 00V\n\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 420  counts\n\r
     Boom angle  = 0.0 degrees\n\r
     True North  = 420  counts\n\r
     Prop Pitch  = 306  mm/rev\n\r
\n\r
d1775\r
D1775\r
\n\r
q\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 420  counts\n\r
     Boom angle  = 177.5 degrees\n\r
     True North  = 168  counts\n\r
     Prop Pitch  = 306  mm/rev\n\r
\n\r
\r
\r
\n\r
W:0005    596    78   252   50 00Y\n\r
W:0005    612    80   253   50 00j\n\r
W:0005    604    79   249   50 00\\n\r
W:0005    596    78   246   50 00V\n\r
W:0005    589    77   246   50 00U\n\r

102: CO2, Site tower, Thu 03-Jan-2002 15:27:47 MST, li7500 calibrations
Here is what I think the Li7500 is reporting (I guess they are using g-moles): 

H2OD - mmol/m3

convert to mass mixing ratio by:

H2OD (mmol/m3) * 0.001 (mol/mmol) * 18 (g_H2O/mol_H2O) / rho.air (kg_air/m3)

CO2D - umol/m3???

convert to ppmV by:

CO2D (mmol/m3) * 1000 (umol/mmol) * 28.97 (g_air/mol_air) * 0.001 (kg_air/g_air)
 / rho.air (kg_air/m3)

A plot of values converted this way is attached.  H2OD is compared to Q and 
CO2D is compared to co2mr.  Everything seems in the right ballpark.

101: Snow_Precip, Site tower, Thu 27-Dec-2001 18:59:25 MST, New temp variable from snow depth gauge
The cr10 that is sampling the snow depth gauge also outputs
a temperature used in the calculation, which we have not been
parsing (and therefore does not appear in the archive).

At 17:25 MST Dec 26 cosmos was rebooted with a new sio_converter
for channel 201 which parses the temperature. 

The new variable was first called Tsnow - which is perhaps a bad 
choice since it isn't a snow temperature.

Later the snow depth variables were renamed and covars rerun back
to Dec 10. (In the field version of the netcdf files the original
names will still exist in files up to Dec 28).

	depth.snow.main	->  Depth.snow
	Tsnow.main	->  Tdepth.snow




100: Temp_RH, Site all stns, Wed 26-Dec-2001 12:13:04 MST, temperature event?
For a few hours before dawn this morning, the bare and sage temperatures were
several degrees colder than at the main tower.  The only quick explanation I
can come up with is that the main tower was in a different air mass than the
other two stations.

99: Radiometers, Site tower, Wed 26-Dec-2001 12:02:22 MST, Rsw.in(main) odd
Rsw.in at the main tower has a lot of extra variation at night compared to sage
and compared to both Rsw.out.  This has been going on for a long time -- perhaps
since the beginning.  Could its fan be bad?

98: , Site all stns, Wed 26-Dec-2001 11:55:35 MST, temp sensitivities
I now suspect temperature sentivities of several sensors:
- CO2 closed-path
- CO2 open-path
- (ASTER?) Barometer
- Snow depth gauge

All should be characterized in some fashion:
- CO2 closed-path by processing cal gas values
- CO2 open-path by comparison to corrected closed-path???
- Barometer through cal chamber check (post cal)
- Snow depth through "stool" check (also post cal in chamber)

97: Snow_Precip, Site tower, Wed 26-Dec-2001 11:52:13 MST, temp sensitivity
I would guess that the snow depth gauge has a temperature sensitivity, though
perhaps it drifts as well.   I just tried:
    plot(dat("T")[,"T.1m"]*2.5+2100-dat("depth"),ylim=c(-200,200))
which looks like it removes some drift, but perhaps gives a signal which isn't
correct either.

Add to the list of one-time tasks:
- put stool under snow gauge
- let it sit for a period during which the temperature changes (several hours)
- remove stool
- log "stool" time in logbook

96: Base, Site all stns, Wed 19-Dec-2001 10:54:00 MST, archive size and disk capacity
Archive file sizes (Kbyte = 1024 byte) per day:

	marigold: 209396
	cosmos:   177885
	s2:        36642
	s3:        36987
        aster:       185

	total:    461095

The /data disk is: 11442156 Kbytes, so we can store 24 days on /data.

Backup disk #2 is 16547932 Kbytes, which holds 35 days.

Backup disk #1 is slightly larger than #2, something like 18 Gbytes.

Note that these sizes don't include any photos.

95: Temp_RH, Site stn2, Tue 18-Dec-2001 12:13:43 MST, TRH fan replaced
The TRH fans were swapped out with low-temp silicon oil.

94: Sonics, Site stn3, Tue 18-Dec-2001 12:06:48 MST, Sage site sonic boom angle
Readings looking toward csat:
Theodolite				90.47
KVH#3 mounted on boom (w.o. decl.)	77.3
Hand-Held compass (12 deg. decl.)	91
DataScope (w.o. decl)			75.3

Declination of site 12 - 12.5 degrees per FAA sectional


Add 180 and convert to degrees to get boom azimuth:

270.47 deg
Add 90 to get Vazimuth: 0.47 deg


Sunangle data:
	Input TIME window desired:
		Starting hour: 12
		Number of hours:   2
		time interval in minutes:  5
		No. of Hours behind Greenwich Time:7


			Sun Angle Data
			     for

		earth location....latitude:   40,40,11
				  longitude: 106,19, 3
		on date: 12/18/1901 

local time,    Z time         Grnwich    earth      local sun position  
hour/minute                  Hour Angle  decl.      azimuth    elev.  
-----------    ------        ----------  -----      ------------------

 12: 0         19.00          105.83    -23.39       179.50    25.94
 12: 5         19.08          107.08    -23.39       180.78    25.94
 12:10         19.17          108.33    -23.39       182.05    25.92
 12:15         19.25          109.58    -23.39       183.33    25.87
 12:20         19.33          110.83    -23.39       184.60    25.81
 12:25         19.42          112.08    -23.39       185.87    25.72
 12:30         19.50          113.33    -23.39       187.14    25.61
 12:35         19.58          114.58    -23.39       188.40    25.48
 12:40         19.67          115.83    -23.39       189.66    25.34
 12:45         19.75          117.08    -23.39       190.91    25.17
 12:50         19.83          118.33    -23.39       192.16    24.98
 12:55         19.92          119.58    -23.39       193.41    24.77

 13: 0         20.00          120.83    -23.39       194.64    24.54
 13: 5         20.08          122.08    -23.39       195.87    24.29
 13:10         20.17          123.33    -23.39       197.09    24.02
 13:15         20.25          124.58    -23.39       198.31    23.73
 13:20         20.33          125.82    -23.39       199.51    23.42
 13:25         20.42          127.07    -23.39       200.71    23.10
 13:30         20.50          128.32    -23.39       201.90    22.75
 13:35         20.58          129.57    -23.39       203.08    22.39
 13:40         20.67          130.82    -23.39       204.24    22.01
 13:45         20.75          132.07    -23.39       205.40    21.61
 13:50         20.83          133.32    -23.39       206.55    21.20
 13:55         20.92          134.57    -23.39       207.69    20.76


93: Logistics, Site tower, Tue 18-Dec-2001 09:09:49 MST, guy wire tension
Guy wire tensions Updated:


	lower	middle	upper
NE	520	650	550
SE	690	620	480
SW	620	450	670
NW	620	490	630

This was in about 6 m/s winds from the SW.

(The forcast is for very strong winds (50 mph?) today, so I thought I should
check.)

Update: 9 Dec:
NE	510	650	550
SE	670	590	480
SW	560	430	660
NW	600	480	600
(within 20 of the 6 Dec readings).  Winds about 4 m/s from the S.


Update: 18 Dec:
	L	M	H
NE	500	600	520
SE	630	550	440
SW	570	400	590
NW	600	475	600






92: Temp_RH, Site tower, Mon 17-Dec-2001 17:31:28 MST, TRH fans replaced
The TRH fans were swapped out with low-temp silicon oil.
91: Radiometers, Site stn2, Mon 17-Dec-2001 17:26:43 MST, 4-Component Radiometer Stn2 moved to Stn3
Q7 installed in place of 4-component.
Soil Sensors remain.
Coefficients for Q7 changed at about 00:20Z 18Dec01
Q7 at 1.97m oriented south on guyed T-post.
90: Eve, Site stn3, Mon 17-Dec-2001 17:24:51 MST, PAM Stn3 / sage site setup
17Dec01
Station Set Up approx. 1km N. of base in Sagebrush.

4-component radiometer system from bare site moved to
this location looking down over a lovely specimin of a sage bush.


Q7		1.96m
CSAT #538	2.35m
	Level X(lateral)=-.015   Y(longitudinal)=-.005
	per electronic level mounted on boom
Krypton #1395
Serializer #6


Krypton installed on AM of 12/18 ~17z
Soil Sensors installed on AM of 12/18 ~17z
	Soil sensors not the best in frozen ground.
	Back-fill of ~1/2 of CS615 because unable to
	insert 'poker' far enough to match length of probles.
	Tsoil installed with minimal angle ending at 5cm.
	HFT, CS615 at 5cm.







89: Radiometers, Site all stns, Thu 13-Dec-2001 10:45:06 MST, calibration files for CR10Xs
Below yare the calibration coefficients used in
each CR10X.
NOTE: Before the installation of stn3 at the "sage" site, stn2 at "bare"
had the 4 comp system. This system was moved to stn3.
The soil sensors and the Q7 stayed at stn2 and a new
set of soil sensors and Q7 along with the 4 comp were
installed at stn3. So if you are looking at data for
stn2 before the installation of stn3, you need to use
the 4 comp coefficients listed in stn3.

TOWER

sensor	SN	     location	calibration
Q7	Q99258	A+	00	9.35
		A-	01	11.55
PYGIN	26416		02	280.1120448
PYGOUT	29260		03	311.5264798
KZIN	940187		04	100.7049345
KZOUT	940186		05	82.98755187
HFT	H983212		06	38.4
PIcase			10	2500
		A0	11	0.00335513
		A1	12	0.278704
		A2	13	1.4317
		A3	14	998.718
PIdome			15	2500
		A0	16	0.00305339
		A1	17	0.268816
		A2	18	8.77845
		A3	19	-2267.84
POcase			20	2500
		A0	21	0.00335427
		A1	22	0.279319
		A2	23	5.19638
		A3	24	-247.625
POdome			25	2500
		A1	26	0.00305418
		A2	27	0.268098
		A3	28	4.06691
		A4	29	403.649
Tsoil	NCAR002	A0	30	-241.477805
		A1	31	231.052081
		A2	32	9.262642
level  N0383 X, A0	33	0.05750311
		A1	34	14.26792000
		A2	35	-0.13395810
		A3	36	-0.20585990
		A4	37	0.0000000
		A5	38	0.0000000
	     Y, A0	39	-0.02092993
		A1	40	14.44724000
		A2	41	0.12732100
		A3	42	-0.39985070
		A4	43	0.0000000
		A5	44	0.0000000
CS615	12	A0	45	-0.187
		A1	46	0.037
		A2	47	0.335
ice on/off 1=on, 0=off	50	1
min bat	(volts)		51	11
X time	(secs)		52	300
outmode	0=a, 1=b	55	0


stn2 at bare site
sensor	SN	     location	calibration	
Q7	Q97035	A+	00	9.26	
		A-	01	11.52	
HFT	H993563		06	38.9	
Tsoil	NCAR001	A0	30	-241.6249	
		A1	31	230.9534	
		A2	32	9.0362	
CS615	13	A0	45	-0.187	
		A1	46	0.037	
		A2	47	0.335	

stn3 at sage site
sensor	SN	     location	calibration   (sensors at stn2 before stn3)
Q7	Q94125	A+	00	8.64			
		A-	01	10.78
PYGIN	31978		02	280.8594299		*
PYGOUT	31976		03	259.6053998		*
KZIN	940185		04	80.25682183		*
KZOUT	970380		05	86.50519031		*
HFT	H993560		06	37.7
PIcase			10	2500			*
		A0	11	0.00335405		*
		A1	12	0.28145			*
		A2	13	3.04961			*
		A3	14	289.824			*
PIdome			15	2500			*
		A0	16	0.0030332		*
		A1	17	0.291681		*
		A2	18	-3.4779			*
		A3	19	870.025			*
POcase			20	2500			*
		A0	21	0.00335116		*
		A1	22	0.285883		*
		A2	23	0.220836		*
		A3	24	756.272			*
POdome			25	2500			*
		A1	26	0.00302872		*
		A2	27	0.297063		*
		A3	28	-5.33875		*
		A4	29	1070.53			*
Tsoil	NCAR003	A0	30	-242.721555
		A1	31	233.953195
		A2	32	8.282468
level  N2018 X, A0	33	-0.3555644		*
		A1	34	11.0103000		*
		A2	35	-0.1015489		*
		A3	36	-0.4473981		*
		A4	37	0.0000000		*
		A5	38	0.0000000		*
	     Y, A0	39	0.5774926		*
		A1	40	10.6420400		*
		A2	41	0.0040844		*
		A3	42	-0.1290564		*
		A4	43	0.0000000		*
		A5	44	0.0000000		*
CS615	8	A0	45	-0.187
		A1	46	0.037
		A2	47	0.335
ice on/off 1=on, 0=off	50	0
min bat	(volts)		51	0
X time	(secs)		52	0
outmode	0=a, 0=b	55	1



88: Power, Site stn2, Wed 12-Dec-2001 08:31:45 MST, Changed Icharge monitor to Iwind/Isolar
Monday, 10Dec01, ~14:30 (20:30Z)

New EVE configuration: fmt0113

Purpose: Change in Battery/Charge Box.
	Added new charging current shunt to monitor PV / Wind
	separately.

New Base Parameters:
	Channel 100 same 'Vbatt'
	Channel 101 now  'Iwind'  was 'Icharge'
	Channel 102 same 'Iload'
	Channel 103 new  'Isolar'

Operations Notes:
1)  There will be 'noise' on Iload when the wind generator kicks
    in, I expect that.  The cause appears to be that the shunt is
    on the 'ground' line and there are other return paths for the
    real charging current, which is through the Iload shunt.
    I need to decide how and whether or not to attack that problem
    (not too keen on having the shunt on the '+12' side and being
    routed to EVE.)
    Result: Noise / false reading on Iload when wind generator is going.
	    Iwind will be lower than the actual charging current.

2)  Still no monitor on the charge divert circuit
    for when the batteries are full and the juice gets sent to the
    dummy load.


EVE Config Declaration:

	# WAS....SINGLE CHARGING CURRENT MONITOR....
	#ANALOG: Ad .5Hz FASTOUT=100
	#Batt-v          0       1       0.026093        0
	#Charge-i        1       10      0.0942          0
	#Load-i          2       100     0.0047          0
	#:

	# NOW....
	ANALOG: Ad .5Hz FASTOUT=100
	Batt-v          0       1       0.026093        0
	Charge-i        1       10      0.0942          0
	Load-i          2       100     0.0047          0
	ChargePV        3       10      0.0942          0
	:



87: Base, Site base, Mon 10-Dec-2001 18:24:44 MST, aster rebooted several times
I just rebooted aster several times because I swapped /backup before remembering
to put $ASTER/projects/FLOSS on it (with all the photos that I promised to 
various visitors today (The Star reporter and the Walden science teachers.)

I think everything is back to the way it should be now.

We're leaving this lovely place.

Bye bye...
86: Radiometers, Site tower, Mon 10-Dec-2001 13:36:09 MST, Change of sw.out at main
Between 11:50 - 12:10 at the main radiation site the down-looking K&Z was 
swapped out
K&Z 940186 and its internal to the housing cable was removed.
K&Z 940181 and its associated cable was installed.

All domes were cleaned, all fans were checked and the Q7 checked for level.
The frost sensors were covered for a few minutes

Ninety minutes later the radiometer data were viewed. The new radiometer shows adata stream comparable to the corresponding data stream at the Bare Eve site.
85: CO2, Site tower, Mon 10-Dec-2001 11:02:03 MST, new licor7500 converter
On Dec 10, 09:36 MST cosmos was rebooted to install a
new sio_converter for the licor 7500 on channel 212.

Up until 09:36 we were using noconv_nl as the sio_converter
for the licor, which does no conversion and passes the
ASCII to the cal routine (licor_7500) to parse.

For efficency sake, were now using this sio_converter:

licor_7500 9 "(Data (Ndx %f)(CO2Raw %f)(CO2D %f)(H2ORaw %f)(H2OD %f)(Temp %f)(Pres
%f)(Aux %f)(Cooler %f))"

The cal routine is then the generic serial_float, though I imagine
we may have to create a special one if we want to convert the
units.

The licor periodically (once a second?) puts out a diagnostic
record, which will be ignored (and not archived).  To see
the diagnostic record in real-time use rserial.
84: Base, Site base, Mon 10-Dec-2001 11:01:16 MST, restarting serial_ingest
To effect a new sio_converter for the pam data on stn2 you only
needed to send a HUP signal to the stn2 ingest process:

            ------------- Ingest tasks -------------
host            user   pid    start exectime process
aster       aster     1216 Dec09 ? 00:00:12 stn2 4 /dev/usb/ttyUSB0


kill -HUP 1216

serial_ingest then restarts another ingest on that serial port
immediately.


83: Radiometers, Site tower, Mon 10-Dec-2001 09:31:30 MST, level sensors temperature error
It is obvious that the level sensors at the main site respond inversely to
temperature -- about 1 angle degree to -0.03 C.  Both X & Y show this behavior.

The little bit of data that I have from the bare site don't show this effect.
82: Base, Site all stns, Sun 09-Dec-2001 18:26:41 MST, rebooted aster
The reason why Xlevel and Ylevel have been bad is that channel config had too
few values for the logger at the bare site (cr10_13, rather than cr10_18).
Since I don't know how to restart ingest for the PAM sites (I tried
ingest_control -r stn2, but this didn't work) and I couldn't reach Gordon,
I simply restarted aster.  This appears to have worked (though I still had
to start cockpit stn2 manually).

Now Xlevel and Ylevel.bare appear to be okay.

The biggest problem with the way we were collecting data is that we dropped
all the Qsoilraw.bare values.  Thus, it will be extra work to revise the 
calibrations for the Qsoil.bare data taken to date.

81: Radiometers, Site stn2, Sun 09-Dec-2001 17:03:53 MST, rads cleaned?
I attempted to clean the bare site radiometers about 1640 using cotton swabs.
Both pygs had a fair amount of dirt, especially near the base of the dome.
I also cleaned some obvious, though small, streaks on the top psp.

Note: Radiometers should be cleaned in daylight (preferably nearly clear),
so that the effect of cleaning can be quantified.  I realized too late that
it would have been better to do this cleaning in more sun.

Also note that it is hard to do clean radiometers (especially psp's) well with
a cotton swab.  We should have some clean wipes here.
 
80: Kryptons, Site stn2, Sun 09-Dec-2001 17:03:17 MST, krypton cleaned
I finally got to the bare site to clean its krypton about 1600.
79: Status, Site all stns, Sun 09-Dec-2001 13:28:38 MST, visitors
Bob Kelly (Univ. of Wyoming) was here for a tour this morning.

Other visitors we've had:

Ed - (the local NWS cooperative observer for the past 25 years) was here for
a tour Friday afternoon.

Todd S- and Pam - from the Wildlife Refuge were here Thursday to teach Larry
and Dean about local plants.

Mark Lanier from the Wildlife Refuge was here Wed? last week to see what we
were up to.

We've (mostly Larry) also fielded reporter phone calls from the Denver Post 
and Colorado Springs Gazette last week.

We're expecting a visit by Matt Landis and other local High School teachers
tomorrow afternoon.

78: Radiometers, Site tower, Sun 09-Dec-2001 11:14:41 MST, rads serviced
I opened up Rsw.out to see if I could see a bad connection, but everything 
looked okay.  It was opened about 9:30 and closed about 11:00 and the ladder
was underneath, so this morning's data are bad.

I also attempted to clean Rsw.out with distilled water, but it immediately 
froze.  There still was ice on it when I left.  The others I simply wiped with
a cotton swab.
77: Kryptons, Site tower, Sun 09-Dec-2001 11:12:03 MST, kryptons cleaned
Larry and I just cleaned all the kryptons, starting at 20m over the last hour.
Of course, we had to pull up the booms at 20 & 10, so sonic data will be bad.

76: Status, Site all stns, Sat 08-Dec-2001 17:29:35 MST, shallow inversion today
We noticed a very shallow inversion this afternoon, where the max temperature
was at 5m.  The temperature differences were about 0.2 C from 0.5m to 5m and
back from 5m to 20m.  I believe the temperature sensors by now.  Later this
evening, it broke down and became completely stable throughout the profile.

75: Thermocouples, Site tower, Sat 08-Dec-2001 16:59:37 MST, replaced last thermocouples
I just put the last 2 good thermocouples in at the lowest 2 levels: 0.20 and
0.35m.  I also removed the dead thermocouples from the top down to 8m.
Some were pretty well battered.

74: Logistics, Site stn2, Sat 08-Dec-2001 16:15:42 MST, staked rad cable
I just put the rad cable up on fence posts.  I'm now declaring this site
critter-proof.  (Ha!)

73: CO2, Site tower, Sat 08-Dec-2001 14:21:45 MST, li7500 up
I mounted OSU's LI-7500 yesterday about 2pm and connected it to cosmos:212.
We are archiving the data in unconverted ASCII until we can figure out a
different format.  

I don't like the mounting for the LiCor's box.  Currently, I have it just
sitting on a breakout box stand.  At the very least, it should be bolted to
this stand....

I also did some CONFUSING cabling for this sensor.  I have connected it to
MARIGOLD's Serial/Analog breakout box, to avoid adding another BOB.  My
primary reason was to avoid running more DC power cable, since the LiCor takes
2 AMPS.  A secondary reason was that we might connect it to marigold later
anyway, if Larry decides to move this sensor to 10m.

We still need to parse the message from this sensor.  I've just put a generic
prep.config entry in for this sensor, commented out.

72: Temp_RH, Site all stns, Sat 08-Dec-2001 13:59:34 MST, TRH 703
We now have in the trailer two "spare" T/RHs.  One is 704 which I just removed
from the tower at 20m due to temperature readings that were 0.2 C low (the 
profile looks MUCH better now).

The other is 703 which Steve had removed from 2m a while back.  Steve said that
the fan was bad.  I took it apart and found that the ribbon cable was cut.  I
spliced in a section of normal cable, but noticed some bare insulation on the
wires directly connected to the fan.  I didn't patch these.  Thus, 703 should
now be considered a working spare, though it should have a better repair job
sometime.


71: Soil, Site tower, Sat 08-Dec-2001 12:41:46 MST, soil samples
About 11:40 I took another core at the main site.  I also took a nearby TRIME 
reading of 10.4%.  This time, I cleared a bare spot, but there were still a lot
of roots (one ~4mm diam.) in the sample, so most of the sample felt like peat.

The was separated into layers.  The top cm was only partly filled, since the
sample was on a small local slope.  I estimate the average depth in this as 
0.5cm.  Also, the bottom wasn't completely filled either.  I estimate the depth
as just 2cm.

I simply weighed the total dirt sample (wet) in each of 4 layers to get the
total bulk density, then created alloquots to determine the soil moisture
after drying.  The corer diameter is about 5.4cm

Depth Cont. Tare Tare+Samp Samp         Ave moist  Comp.DryWeight DryDens
0-1.5   -   28.9  53.0     24.1
1.5-4.5 B   26.9  69.6     42.7
4.5-7.5 C   27.0 119.5     92.5
7.5-9.5 D   27.0  95.8     68.8

For the top 4.5 cm, bulk density is g/cm^3

Depth  Alloquot  Wet   Dry   Water
0-1.5   A1	 13.3   2.9
        A2	 10.9   2.4
1.5-4.5 B1  	 27.0  12.2   
        B2       15.4   9.7
4.5-7.5 C1       47.9  43.1 (slightly spilled before dry weighing.)
        C3       33.8  30.5
7.5-9.5 D1       33.8  31.7
        A3       34.7  32.8

Since this sample was heavily influenced by the grass, I'll get another
tower sample today.



70: Status, Site all stns, Sat 08-Dec-2001 10:12:55 MST, serial numbers and heights
These are the serial numbers for sensors that I know of:

Main:	TRH	Prop	CSAT	KH20	Baro	Logger
20	704	0007	0540	1394
15	701	0006
10	201	0003	0537	1393
5	001	0002	
2	004	0001	0536	
1	702	0004
0.5	005
Bare:
2	007		0539	1392	1002	1

Note that I'm just about to change the TRH at 20m to SN 104. (Done 10:55.)

Here are the heights of these sensors:

TRH
20	20.40+0.31-0.04-0.57 = 20.10
15	16.72-0.72-0.04-0.57 = 15.39
10	11.20-0.67-0.04-0.57 = 9.92
5	5.78+0.31-0.04-0.57 = 5.48
2	2.10+0.46-0.04-0.57 = 1.95
1	2.10-0.55-0.04-0.57 = 0.94
0.5	2.10-1.02-0.04-0.57 = 0.47

Prop
20	20.40-0.84-0.06+0.45 = 19.95
15	14.88-0.01+0.45 = 15.32
10	9.46+0.05-0.06+0.45 = 9.90
5	5.78-0.86-0.06+0.45 = 5.31
2	2.10-0.51-0.06+0.45 = 1.98
1	2.10-1.57-0.06+0.45 = 0.92

CSAT
20	20.40-0.89-0.06+0.05 = 19.50
10	11.20-0.89-0.06+0.05 = 10.30
2	2.10+0.04+0.05 = 2.19

KH20 relative to CSAT: 20cm down, 35cm back
LI6251 inlet frit relative to CSAT: 10cm down, 33cm back
LI7500 center of path relative to CSAT: 40cm down, 67cm back





		20.40-0.84-0.06+0.45
69: Temp_RH, Site tower, Sat 08-Dec-2001 10:11:28 MST, T/RH swapped ~11/30
Shortly before he left, either 11/30 or 12/1, Steve Semmer swapped T/RHs at
2m.  004 replaced 703.

68: Soil, Site tower, Sat 08-Dec-2001 09:11:11 MST, soil samples
About 2:00 I took a core at the main site.  I also took a nearby TRIME readings
of 5.7% and 6.9%.

I still had trouble unscrewing the corer when I got it back to the trailer, but
eventually got separated it out in layers.  However, the top 2cm ended up being
above-ground grass, so the core was not as deep as it should have been.  Also,
the first 3cm of ground, and even some of the next 3cm was mostly roots (like
peat).

I simply weighed the total dirt sample (wet) in each of 4 layers to get the
total bulk density, then created alloquots to determine the soil moisture
after drying.  The corer diameter is about 5.4cm

Depth Cont. Tare Tare+Samp Samp         Ave moist  Comp.DryWeight DryDens
0-3   B     26.9  59.7      32.8 (peat)    (39%)   23.6		  0.34
3-6   C     27.0 116.1      89.1 (clumps)  (11%)   80.3		  1.17
6-8.7 D     27.0 129.1     102.1 (broken)  (8%)    94.5		  1.65

For the top 6 cm, bulk density is 0.76 g/cm^3

Depth  Alloquot  Wet   Dry   Water
0-3    B1        11.8   8.1   3.7 (31.4%)
       B2        12.3   6.6   5.7 (46.3%)
3-6    C1        32.9  29.4   3.5 (10.6%)
       C3        32.0  28.1   3.9 (12.2%)
6-8+   D1        40.8  37.7   3.1 (7.6%)
       A3        36.1  33.3   2.8 (7.8%) spilled during second weighing, but 
I think I recovered most of it)

Since this sample was heavily influenced by the grass, I'll get another
tower sample today.

67: GPS, Site base, Sat 08-Dec-2001 09:09:45 MST, gps died again
The GPS was dead this morning.  I'm going to guess that the power supply doesn't
like the cold of the "knee-hole", since it was quite cold last night (-13 now).
I got it going, but will move the power supply out soon.

66: Sonics, Site stn2, Fri 07-Dec-2001 17:52:04 MST, sonic interruptions
I put my mitten in the sonic path for 2 seconds to see if the serializer heater
flag would change (from 256).  It didn't.  Sorry for the data interruption.


65: CO2, Site tower, Thu 06-Dec-2001 16:50:19 MST, flow rates checked
I just did the flow rate checks that Tony suggested.  Each cal gas was turned
on for about 30 seconds at about 16:25.  Note that I was standing near the
sonic during this period, so all fluxes at 2m are suspect.

nominal flow rate is ~5.5 lpm
flow rate with cylinder 1 is ~-2 lpm
flow rate with cylinder 3 is ~-2.5 lpm

These readings were so close to what Tony intended that I didn't change
anything.  

But we still don't know why his gas is being used up relatively fast.  I can
think of several possibilities:
- The flow rate out of the pressurized cylinders is more than through the
scintered inlet.  I could check this by metering the analyzer outflow.
- There is a leak in the cal gas line.  This strikes me as unlikely since both
cylinders are going down at about the same rate.
 

64: Thermocouples, Site tower, Thu 06-Dec-2001 14:50:05 MST, Lower thermocouples: changes to cr21x
The cr21x is now working in high resolution binary mode, as of 14:35 MST,
Dec 6. The cal routine is camp_hr_bin.  To process the earlier data,
one would have to use the serial_float routine. If it is important
we will have to create a new ops period.

The binary data mode fixes the problem of the broken records
coming out of the cr21x - it was putting a carriage return after 
8 fields in the ASCII output.
63: Soil, Site tower, Thu 06-Dec-2001 13:36:51 MST, soil readings
This page summarizes the TRIME readings:

	main		bare
	time	reading	time	reading
Dec 4	1015	 7.1%	1615	 4.7%
Dec 5	1230     7.5%	1310	11.4%
Dec 6	1330	 8.8%	1030	12.3%
Dec 7   1400     6.3%   1700     6.9%
Dec 8   1140    10.4%   1600     7.3%
Dec 9   1245     8.0%   1630     9.1%
Dec 10  1230     5.6%	1500	 9.4,9.7,9.9%

I've plotted these values with the 615 readings.  Obviously, the comparison
is pretty bad -- probably this is a spatial sampling issue.









62: Soil, Site stn2, Thu 06-Dec-2001 12:34:56 MST, soil samples
About 10:30 I took a core at the bare site.  I also took a nearby TRIME reading
of 12.3%.

I had trouble unscrewing the corer when I got it back to the trailer, so I
simply ground dirt out, making the profile sampling a bit crude.

I simply weighed the total dirt sample (wet) in each of 4 layers to get the
total bulk density, then created alloquots to determine the soil moisture
after drying.  The corer diameter is about 5.4cm

Depth Cont. Tare Tare+Samp Samp         Ave moist  Comp.DryWeight DryDens
0-2   -     28.9 78.1       49.2 (damp) 12.7%       43.7	  0.95 
2-5   B     26.9 139.9     113.0 (wet)  12.3%	   100.6          1.46
5-8   C     27.0 125.7      98.7 (wet)  8.1%        91.3          1.33
8-10.7 D    27.0 108.4      81.4 (dry)  8.6%        75.0          1.21

For the top 5 cm, bulk density is 1.26 g/cm^3

Depth  Alloquot  Wet   Dry   Water
0-2    A1        18.8  16.5  2.3 (12.2%)
       A2        19.6  17.4  2.2 (11.2%)
2-5    B1        23.8  20.8  3.0 (12.6%)
       B2        24.7  21.7  3.0 (12.1%)
5-8    C1        34.8  31.7  3.1 (8.9%)
       C3        36.4  33.7  2.7 (7.4%)(I added one of the 2 rocks)
8-10+  D1        30.9  28.0  2.9 (9.4%)
       A3        30.6  28.2  2.4 (7.8%)

Tomorrow, I'll go for a sample at the tower.


61: Soil, Site tower, Thu 06-Dec-2001 12:27:02 MST, soil sampling
Procedure for soil samples during FLOSS01

At both sites, somewhere in the vicinity of the radiation stand/soil sensors, 
take periodic soil cores for Bulk Density and for Soil Moisture.
Bulk density need only be done a few times at each site, but it would be good to do
  soil moisture periodically.
I tried out the soil corer during set-up and found it to be a suitable method
for both Bulk Density and Soil Moisture
After using the corer clean it and wipe the threed with a WD40 tissue.
Note that the barrel need not be screwed in too tightly

Assemble the core with two 1cm sleeves then one 3cm sleeve and then the remaining
sleeves.
 This will enable the top 5 cm to be separated from the core.
At the sample location, clean away the snow and loose dead grass.
Using the slide hammer pound the core into the soil until it bottoms out.
Work the corer back and foreward, left and right until it is loose.
Lift the corer out and carry it back to the base.
Use the plastic box in the tool box as a work area.
Unscrew the barrel of the corer and push out the sleeves containing the soil core.
Using the flat bladed tool separate the core at the 5 cm level.

To determine bulk density (and soil moisture):
Place entire 5cm core in one of the larger Al containers.
Weigh, setting the tare with a similar Al container.
Cook for two days in the oven at 110 C.
Reweigh.
From volume of the core (5cm x pi r^2 cm^2) and dry weight calculate bulk density.
From fresh weight and dry weight calculate the gravimetric water fraction.
          (fresh wt  - dry wt)/dry wt = gravimetric water fraction.

To determine soil moisture only (but more accurately).
After separating the top 5 cm of the soil core break up the soil mass and
homogenize it.
In three of the small Al containers (labelled A1,A2,A3, B1,B2,B3, C1,C2,C3) load ~
30g of soil.
Weigh, using a similar container as tare.
Cook for one day in the oven set at 110 C.
Reweigh.
From fresh weight and dry weight calculate the gravimetric water fraction.
          (fresh wt  - dry wt)/dry wt = gravimetric water fraction.

60: CO2, Site tower, Thu 06-Dec-2001 12:26:23 MST, CO2 installation
The Licor aspirated CO2 system was installed on 30 Nov''01
The configuration is shown in ...../doc/co2licor.idr (to be copied from
~/deploymnets/floss02/co2licor.idr)

lico2.2m cosmos 102
li.2m    cosmos 103

         Channel 0   Channel 1
Ambient      off         off
Lo-cal       on          Off
Hi-cal       off         on

Sequence of 2 minutes Lo-cal, 2 minutes of Hi-cal, 4 Hr 56 minutes of Ambient,
repeating every 5 hrs
was initiated.

The system was tested on 1 dec''01.
Lo-cal and Hi-cal were activated and a Zero can period was run.
The Flows were checked with a floating ball meter.
5 lpm intake during ambient flow.
    Adjusted using bypass valve on pump suction manifold.
    
~2 lpm backflow during both Lo-cal and Hi-cal periods.
    Adjusted using cross handle, very sensitive, for lo-cal
         and using knob, for hi-cal

Set a default calibration factor of 180 ppmv/volt to generate an approximately
correct mixing ratio
Keeping track of cal gas usage

59: Radiometers, Site all stns, Thu 06-Dec-2001 12:25:28 MST, radiometers and soil sensors
Radiometers and soil sensors

Radiation stands were set up at the Main site and the Bare site on 28 Dec''01

The radiometers were installed at Bare on 29 Nov''01
The radiometers and snow sensor were installed at Main on 30 Nov''01

The soil sensors were installed at Main on 30 Nov''01
The soil sensors were installed at Bare on 30 Nov''01

The radiometers were leveled and fan function were checked at Main on 01 Dec''01
The radiometers were leveled and fan function were checked at Bareon 01 Dec''01

The complement of sensors is shown in ...../doc/radioms.idr (to be copied from
~/deploymnets/floss02/radioms.idr)

58: CO2, Site tower, Thu 06-Dec-2001 12:23:08 MST, cal gases
On 8 Nov''01 six cylinders of synthetic air/CO2 were cross calibrated with a CMDL
standard at CUFF.

Cal1 AK-170118 317 ppmv

Cal2 220K-0017 315 ppm

Cal3 LK58629   484 ppmv

Cal4 LK285067  495 ppmv

Cal5 291120Y   493 ppmv

Cal6 220K-0002 324 ppmv

Cal5 and Cal6 remain at FLAB
Cal1, Cal2 , the Lo-cal cylinders, and Cal3, Cal4. the Hi-cal cylinders are at the
tower base.
Cal1=Lo-cal and Cal3=Hi-cal were connected Sat 01 Dec''01
Cal2=Lo-cal and Cal4=Hi-cal are in reserve, laying next to the active cylinders.

57: Power, Site stn2, Thu 06-Dec-2001 12:21:22 MST, changed wind gen pot again
I now have the pot all the way clockwise.  This should be the high voltage 
setting.  The prop was much happier and went from spinning slowly to quite fast.
I note that Icharge went up to 14A!  Hopefully, this is a good setting to leave
it at.

56: Radiometers, Site tower, Thu 06-Dec-2001 09:39:42 MST, checked Rsw.out
We noticed (belatedly) last night that Rsw.out.main has been bad.  I reseated
cables this morning about 9:20 and did two tests about 9:30- 20 seconds with
my hand underneath followed by 20 seconds with the guy tension meter wood box
underneath.  I'll look at these data soon.

With my hand, the value was 22 W/m2, with the box 18 W/m2, uncovered about
70 W/m2.  Thus, it appears to be working now.  I hope it stays that way.

55: Logistics, Site tower, Thu 06-Dec-2001 09:34:14 MST, guy wire tension
I just measured guy wire tension:

	lower	middle	upper
NE	520	650	550
SE	690	620	480
SW	620	450	670
NW	620	490	630

This was in about 6 m/s winds from the SW.

(The forcast is for very strong winds (50 mph?) today, so I thought I should
check.)

Update: 9 Dec:
NE	510	650	550
SE	670	590	480
SW	560	430	660
NW	600	480	600
(within 20 of the 6 Dec readings).  Winds about 4 m/s from the S.


54: Thermocouples, Site tower, Wed 05-Dec-2001 13:57:46 MST, TC data from 21X
The data from the 21X is somewhat scrambled.  Every other sample from the
thermocouples is 0.  For the samples that the thermocouple is 0, the reference
appears to be a real value.  The alternate samples appear to be good, with
a slowly varying reference and real thermocouple values.

Either we're not parsing the 21x data properly, or it is being driven too fast.

53: Snow_Precip, Site tower, Wed 05-Dec-2001 13:15:50 MST, precip gauge okay
The tips from the rain gauge during my calibration at about 12:35 were:
5, 7, 7, 7, 8, 7, 2 = 43 tips total
This sounds like a reasonable number.  The CASES97 logbook says that we would
expect 44!

52: Soil, Site stn2, Wed 05-Dec-2001 13:13:29 MST, soil Q
Sampled one location: 11.4%

also shot 3 photos.

Also changed wind gen pot to full counterclockwise.

P.S. There was no charging the following night even with 10 m/s winds, so
this pot adjustment was bad.

51: Status, Site tower, Wed 05-Dec-2001 12:54:23 MST, site visit
I spent the last 20 minutes at the main tower for daily stuff:
- dumped water in rain gauge (appeared to work!)
- shot 3 photos
- measured TRIME moisture 6.9 -> 7.4%

Now off to bare site for TRIME stuff and change wind gen pot.

50: CO2, Site tower, Wed 05-Dec-2001 12:11:51 MST, co2 calibration
I just looked at the CO2 calibration on 5 Dec 02:00.  From an eyeball average
of values from prep, using a default gain of 190 and bias of 0:
cyl #1 - 317 => 334
cyl #3 - 484 => 472

Thus, the gain would be: 190 * (484-317)/(472-334) = 230
and the offset would be: 317 - 334*230/190 = -87

I'll go ahead and put these in, though we expect to generate a time-varying
calibration eventually.

49: Snow_Precip, Site tower, Tue 04-Dec-2001 23:16:03 MST, rainr
This afternoon (~4:00), I reseated the connectors for the rain gauge.  I had
the idea that it wasn't completely seated at the CR10 side.  Now, I see a single
tip which should have been due to the 1/4" of fluffy snow we just had.  I might
have been tempted to try another calibration, but I didn't have any water left.
I've refilled it now and will try again tomorrow.

48: , Site , Tue 04-Dec-2001 22:31:47 MST, new snow
We had some flurries earlier this evening, but we've just started getting big 
flakes now.
47: Status, Site base, Tue 04-Dec-2001 21:34:38 MST, second mouse trapped
After seeing it scurrying across the back of the trailer today, I moved the
trap tonight.  It was filled 20 minutes later.  10 seconds of silence, everyone.

46: Snow_Precip, Site tower, Tue 04-Dec-2001 16:54:01 MST, added stool!
In a further attempt to check the snow depth gauge, I just placed the small
stool underneath it for about 5 minutes.  The stool height is 11 7/8".

The change was obvious even in cockpit.
The numbers are:
2056-1764=292 (should be 300)
45: Thermocouples, Site tower, Tue 04-Dec-2001 16:52:07 MST, 5m thermocouple fixed
I just replaced the 5m thermo., which brought it back to life.
44: Barometer, Site all stns, Tue 04-Dec-2001 15:49:34 MST, Pressure error
I've been puzzled by the pressure difference between main and bare.  The height
difference only justifies about 0.3mb (bare lower).  (The difference in height
from the topos was about 13' = 4m at ~11m/mb.)  So, I just made a difference
plot and noticed that it varied like temperature by 2mb over 15C.  A 
scatterplot of the pressure difference vs. temperature (see attached) has a 
perfect correlation.  This sounds like the same problem we saw in OASIS98.  
We should get rid of the non-compensated sensor.
43: Soil, Site all stns, Tue 04-Dec-2001 14:39:08 MST, soil moisture
I made some hand-held TRIME measurements near the soil probes at the main tower
today.  At two sites I got: 7.3% and 6.9%.  Note that the first readings were
both 6.9%.  The first site showed increasing values over the first minute.

These were done about 10:15 this morning.
 
Later in the day, (4:15), took readings at bare site: 4.2%, 4.7%, and 5.0%.
I think that the 4.7 was about like the 615 probe.


42: Adams, Site tower, Tue 04-Dec-2001 14:01:35 MST, marigold down
marigold was down from 19:45 to 22:28 today (when Gordon called me to tell
me it was down!)

This was during the time that Massman was setting up -- perhaps something was
bumped.  (He was getting power from the box that the Breezecom is in.)

Sorry for not noticing this...
41: Thermocouples, Site tower, Tue 04-Dec-2001 12:27:15 MST, Sean's thermocouple comments
Here's my summary of what I did there:

---------------------------------------------------------------
The FLOSS thermocouple setup was between Nov 30 - Dec 2, 2001
near Walden, Colorado.

Layout of the thermocouples on the FLOSS Main Tower.

tc level data logger    channel
-------- -----------    ----------
 20.0m        23X       channel 12
 18.5m        23X       channel 11
 17.0m        23X       channel 10
 15.5m        23X       channel 9
 14.0m        23X       channel 8
 12.5m        23X       channel 7
 11.0m        23X       channel 6
  9.5m        23X       channel 5
  8.0m        23X       channel 4
  6.5m        23X       channel 3
  5.0m        23X       channel 2
  3.5m        23X       channel 1

  8.0m        21X       channel 8
  5.75m       21X       channel 7
  4.25m       21X       channel 6
  2.75m       21X       channel 5
  2.0m        21X       channel 4
  0.5m        21X       channel 3
  0.35m       21X       channel 2
  0.2m        21X       channel 1


Notes:

1. The 23X data logger is at about 9m.

2. The 21X data logger is on the ground beneath the stairs.

3. The reference temp sensor from the 21X data logger appeared
   to be noisy (need to talk to Steve Semmer about this).

4. Borrowed PVC fitting for the 21X Campbell Box from Larry/Dean.
   (need to get them a replacement plus some putty-type stuff).

5. Three extra thermocouples are in a wooden red box in the ATD trailer.

6. Stored the following at the ATD trailer:  (1) black plastic box
   w/ thermocouple wire and 21X/23X manuals; (2) red wooden box with
   extra thermocouples in them; (3) plastic tupperware box useful to
   transport thermocouples up/down the tower.

7. On Saturday Dec 2 installed 4 thermocouples from 14-18.5m.  The
   following morning 2 of these thermocouples were broken.

8. When I left Sunday afternoon there was one broken thermocouple
   (0.35m)...I think all the others were working.  The wires breaking
   are going to be a problem...possibly due to the low temperatures?

9. These are E-type (Chromel/Constantan) thermocouples with a 
   diameter of 0.0254 mm (1/1000 of an inch).  Perhaps larger diameter
   wires would be more tolerable of the cold...though we may lose
   some response time.

---------------------------------------------------------------

Also, here's an example of an error I see with the 23X data logger (i
did this while I had the logger here in my office):

104 1 23.934 23.381 22.633 -99999  -99999  -99999  -99999  -99999  -99999  -99999
-99999  22.536 22.585
104 1 23.934 23.381 22.649 -99999  -99999  -99999  -99999  -99999  -99999  -99999
-99999  22.472 22.574
104 1 23.934 23.386 22.639 -99999  -99999  -99999  -99999  -99999  -99999  -99999
-99999  22.483 22553  <---- error here
104 1 23.934 23.391 22.606 -99999  -99999  -99999  -99999  -99999  -99999  -99999
-99999  22.52 22.558
104 1 23.934 23.391 22.617 -99999  -99999  -99999  -99999  -99999  -99999  -99999
-99999  22.488 22.531

where it listed the channel 13 value as "22553" it should be 22.553.
Also, I should have mentioned in my previous e-mail that the ref temp
of the data loggers is also recorded (in channel "0").  I also forgot
to mention about the open channels on the 21X data logger not
outputting "-99999"...instead they would output some random number
(which could possibly look real).

Is it possible to get a small 1-hz data file with some of the
thermocouple data in it?  If this is a lot of trouble don't worry
about it...but I'm curious to see what is happening with the 21X
data...
40: Snow_Precip, Site tower, Tue 04-Dec-2001 11:40:19 MST, depth gauge check
I also put some 2x4 boards under the depth gauge from about 10:30-10:40.  First,
about 1 min with 2" (1.5") thickness, then about 9 min with 4" (3.5") thickness.
Cockpit didn't show much change -- I'll look at data now.

The data appear to be:
2070 - pre test
1999 - 2x4 thin     dh = 71mm (should be 38 + grass bed -- 33?)
1962 - 2x4 thick    dh = 37mm = (should be 51)
2062 - post test    dh = 100mm = (should be 51 + grass bed)

These aren't quite right.  Also, I'm puzzled by the timing, since the "1999" values were before I did anything.

I'll try a slightly different test later.
39: Snow_Precip, Site tower, Tue 04-Dec-2001 11:38:20 MST, rain gauge cal
About 10:30, I dumped a whole bunch of water into the rain gauge, but didn't 
see any response on the cockpit display.  I'll look at the raw data to see
if I should have seen it.

I've looked at the data -- there were no non-zero values.  Rain gauge data isn't
getting in.

38: CO2, Site tower, Tue 04-Dec-2001 11:16:59 MST, cal gas values
calibrations from Niwot Ridge site:

cal 1 317
cal 2 315
cal 3 484
cal 4 495
cal 5 493
cal 6 324 ppmV

We are currently using cylinders 1 and 3.

Bill Massman and John ? are now sniffing these as well.

37: Status, Site base, Mon 03-Dec-2001 18:04:08 MST, covars restarted
I just added X&Ylevels to covars and restarted.

36: Communications, Site tower, Mon 03-Dec-2001 17:07:03 MST, raised breezecom antenna
Tried raising the antenna a bit more on a round pole and aligned to the base
antenna.  Made the number of errors worse -- 47% vs. 42%.

35: Power, Site stn2, Mon 03-Dec-2001 16:39:30 MST, wind generator pot turned
I just changed the pot on the wind generator.  Hopefully, this will improve the amount of power we get out of it.  I  the pot was almost completely clockwise, which would have been about 17.8 V.  I changed it to the middle, which should be about 15.8 V.  We'll see tomorrow if this helped (if the winds cooperate,
they just died).

P.S. winds overnight were 3-4 m/s, and Icharge was 0.  I think that this is an
inconclusive test.


34: Adams, Site tower, Mon 03-Dec-2001 15:55:08 MST, unknown crash of marigold and cosmos
For some reason, both marigold and cosmos crashed.  They both came back up with a "mxreset" apiece.  Were Larry & Dean playing with power???

33: CO2, Site tower, Mon 03-Dec-2001 13:15:09 MST, cal gas amounts
We'll start to keep track of cal gas amounts here:

	Cal 1(Lo)	Cal 3(Hi)
01 Dec	20/1750		20/1630
03 Dec	 ?/1710		 ?/1610
04 Dec  22/1610         19/1595 (~10:30; before Massman)
04 Dec  22/1590         20/1480 (~16:30; well after Massman)
06 Dec  22/1400         20/1410
07 Dec  22/1360         20/1400
08 Dec  22/1250         20/1330
09 Dec  22/1200         20/1300
10 Dec  22/1120         20/1290








32: , Site all stns, Mon 03-Dec-2001 10:51:29 MST, task list
Here are the tasks I can think of:

Daily:
- check_aster
- daily status to logbook
- snow photograph
- trime soil moisture reading
- read cal gas gauges

End of Walden visits:
- swap backup disk
31: Status, Site stn2, Mon 03-Dec-2001 10:12:42 MST, misc
I just measured the sonic height as:
79" + 2.5" + 1.2" = 82.7" = 2.10m

I also took several photos from NW-W-SE at about 10:00

I also just measured power.  Solar, battery about 14.16V.  Wind generator 
by itself (disconnected) about 16.86V.  When the wind generator is plugged in,
it runs significantly slower.  This doesn't seem right.  When the battery is
low (and solar down), would the generator run even slower?

30: Thermocouples, Site tower, Mon 03-Dec-2001 09:34:07 MST, last value (20m)
Gordon had noted that if the last 23x thermocouple is missing, that the reported
value is -9, -99, -999, -9999 or -99999, which looked like some characters were
being clipped.  This begged the question of whether characters were being
dropped from normal readings.

I've just looked at a few minutes of values using rserial and don't see any
characters being dropped from normal readings.

29: Status, Site all stns, Mon 03-Dec-2001 08:46:28 MST, daily status
Daily status:

Weather: almost completely clear (some clouds over mountains), snow disappearing
(10-20% bare?).  I'll try to take tower photos today.

cosmos-up
marigold-up (goldenrod hasn't crashed since replaced)
stn2-up
aster-gives RPC error

main site:
props    all okay
T/RH     something has been weird at 15 and/or 20m.  Maybe I'll just swap these.
Tc       5.75m died.  Others okay?  I'll have to make a daily plot that adds in
         the reference.  The 21x values appear to be somewhat noisy.
sonics   okay, though more spikes than I'd like (20-30 overnight at all levels)
kryptons okay, the 2m sensors have had spikes (large variance) in the data.
         I'll clean these soon, though the spikes are probably from blowing snow
licorCO2 okay, cal gases are cycling
rad	 okay I guess.  Haven't looked at daily plots yet.
soil	 okay I guess.  "
icedet   no action (which is good)
levels   okay
baro     okay
precip   okay?  reading zero.  I should do a calibration.
snowdepth okay?  may be seeing blowing grasses
power	 no problems

bare site:
T/RH	 okay?  The fan acted strange last night, perhaps due to power
sonic	 good.  The fewest spikes (perhaps the least blowing snow?)
krypton  okay.  weird values on cockpit probably from my work last night
heater   false high readings.  need to investigate this.
rad      okay.  need to look at daily plots
soil     okay.  "
levels   weird excursions.  need to check this
baro     okay
gps      okay
power    mostly only charging during day -- problem with wind generator?

need to rabbit-proof all cables!




28: Sonics, Site stn2, Sun 02-Dec-2001 17:39:17 MST, bare sonic moved down
About 1/2 hour ago, I moved the sonic down to just above the T/RH crossarm
level on the bare station.  (You should be able to tell when, because I had 
to disconnect the krypton.)  The height of this sensor will have to be 
remeasured.

I also noticed that the T/RH fan went fast/slow/off -- something may be weird
with power.

27: Power, Site base, Sun 02-Dec-2001 16:32:07 MST, Controlling UPS

To run UPS program:

	su -
	cd /usr/lib/powerchute
	./xpowerchute


Under Configuration->Event Actions
  UPS On Battery
     configured to send email to maclean, oncley and semmer
  Low Battery Condition
     configured to send email to maclean, oncley and semmer
     will shut down server after 30 seconds
	

26: Communications, Site all stns, Sun 02-Dec-2001 14:47:55 MST, data network configuration
Data Network

BCAP: Breezecom Access point in base trailer
BCSA: Breezecom Station Adaptors at tower and in OSU box

aster: NCAR base system, Dell Latitude, linux
tiny:  OSU linux system
ambra: Windows 486 running Breezecom DS.11 diagnostic software

****************** 192.168.10.0 *************************************

aster -- 5port --  BCAP ---RF---  BCSA --cable-- 5port ----- cosmos
ambra -- hub        |             tower          hub   ----- marigold
tiny  --            |                                  ----- web cam
                    RF
                    |
                    |-------RF--  BCSA --cable-- 4port ----  cr5000
		                  OSU            hub   ----  lunch

*********************************************************************

These addresses are assigned on the Data Network:

  192.168.10.1	aster
  192.168.10.10-20  allocated dynamically by DHCP server on aster
		    Breezecom radios and Ambra PC

NCAR Data Acq Systems
  192.168.10.30	cosmos	
  192.168.10.31	cosmos-bp	
  192.168.10.32	goldenrod	
  192.168.10.33	marigold	
  192.168.10.34	marigold-bp	
  192.168.10.35	arnica	

OSU PCs
  192.168.10.50 lunch
  192.168.10.51 cr5000




25: Status, Site all stns, Sun 02-Dec-2001 08:33:21 MST, Dec 2 morning status
Weather:
  light snow last night, windy.  Small 8" drifts on refuge road.


Data system:
  up.  Marigold/arnica made it through the night!

Sonics:
  
  some non-zero diag values on all sonics. A few values >=16 which
    indicate a lost sample.
  all sonics also show a few scattered non-zero flags

  heater value at bare site always on

Kryptons: OK
  
TRH: OK

Props: OK

Thermocouples:  Sean had 4 deployed last night.  2 of the 4 are no
  longer reporting, assumed broken.

Radiation: values appear to be in acceptable range - did not do
  a thorough check.  
  Xlevel.bare and Ylevel.bare show some values around 10.

PAM power system:  still some question about whether wind generator
  voltage is getting to charging system.
  Current values: 
	Vbatt: 12.4 V
 	Icharge: 1.5 - 2.6 A  (sun through thin clouds, 7 m/s winds)
	Iload: 1.7 A
	

24: Sonics, Site tower, Sat 01-Dec-2001 17:47:23 MST, increased sampling rate on 20m sonic
data_stats showed that the 20 meter sonic was reporting at 20hz
instead of 60.

I changed it with rserial at 17:45 MST Dec 1,
using the campbell T and Ae commands:

  rserial marigold 201
  ... hex output
  enter "T" to put it in terminal mode
  ctrl-D to get out of rserial

Run rserial in ascii mode:

  rserial -A marigold 201
  ??  display status
  Ae  set it to 60hz
  D   back in data mode
  binary gibberish
  ctrl-D





23: CO2, Site tower, Sat 01-Dec-2001 17:19:58 MST, licor CO2 cal gas sequence
The co2 calibration is controlled by the ttl program in channel config:

# Tony wants a 2 minute calibration every 5 hours
#   17880 secs = 4 hours, 58 minutes
#
# adam          datel_ttl_output unit chan start        on       off
#
cosmos          datel_ttl_output 0 0    00:00:00        120     17880
cosmos          datel_ttl_output 0 1    00:02:00        120     17880

TTL channel 0 controls the low cal gas valve, channel 1 controls
the high cal gas valve.

The low cal gas is 317 ppmV CO2, the high is 484 ppmV.

After cosmos is booted, it performs its first calibration at the
next multiple of 5 hours from 00:00 GMT of the current day, which
is at 05:00, 10:00, 15:00, 20:00 GMT or 01:00 the next day.

The sequence then repeats every 5 hours.

So assuming cosmos was booted at 17:00 GMT, the following sequence
would occur:
	20:00	low cal gas
	20:02	high cal gas
	20:04	ambient air until 01:00 next day.

The next day the cal times will be an hour later:
	01:00	same cal sequence
	06:00
	etc

Cosmos was last booted yesterday, Nov 30.
Tony connected the licor this morning, Dec 1.

I verified in the high rate co2mr.2m data that the calibration 
sequence was done at 21:00 GMT today.





22: Status, Site all stns, Sat 01-Dec-2001 16:23:38 MST, Operational!

All NCAR sensors deployed and reporting data, however quality
checks have been minimal.

Prop angles have been entered, so prop wind directions should be
correct.

Licor co2 is deployed.  See other entry about calibration gas sequence.

Sean Burns has installed 4 (out of 20) thermocouples on the tower.
They are at 14, 15.5, 17, and 18.5 meters.

Kurt, SteveS and Tony left at about 3:00 this afternoon.  SteveO due
back tomorrow.


21: Sonics, Site tower, Sat 01-Dec-2001 12:51:51 MST, Bare site sonic boom angle
Theodolite reading:
 90 deg 17 min 35  sec

Add 180 and convert to degrees to get boom azimuth:

270.293 deg

Add 90 to get Vazimuth: 0.293 deg

Sunangle data:

			Sun Angle Data
			     for

		earth location....latitude:   40,39,56
				  longitude: 106,19,11
		on date: 12/ 1/2001 

local time,    Z time         Grnwich    earth      local sun position  
hour/minute                  Hour Angle  decl.      azimuth    elev.  
-----------    ------        ----------  -----      ------------------

 12: 0         19.00          107.70    -21.90   181 26'23"   27 25'12"
 12: 5         19.08          108.95    -21.90   182 45' 0"   27 23'24"
 12:10         19.17          110.20    -21.90   184  3' 0"   27 19'47"
 12:15         19.25          111.45    -21.90   185 20'59"   27 15' 0"
 12:20         19.33          112.70    -21.90   186 39' 0"   27  8'59"
 12:25         19.42          113.95    -21.90   187 56'59"   27  1'48"
 12:30         19.50          115.20    -21.90   189 14'24"   26 53'24"
 12:35         19.58          116.45    -21.90   190 31'48"   26 43'11"
 12:40         19.67          117.70    -21.91   191 48'36"   26 32'23"
 12:45         19.75          118.95    -21.91   193  4'48"   26 20'23"
 12:50         19.83          120.20    -21.91   194 20'59"   26  6'35"
 12:55         19.92          121.45    -21.91   195 36'36"   25 51'35"

 13: 0         20.00          122.70    -21.91   196 51'36"   25 36' 0"
 13: 5         20.08          123.95    -21.91   198  6'36"   25 18'35"
 13:10         20.17          125.20    -21.91   199 20'24"   25  0'36"
 13:15         20.25          126.45    -21.91   200 34'11"   24 40'47"
 13:20         20.33          127.70    -21.91   201 47'23"   24 20'23"
 13:25         20.42          128.95    -21.91   202 59'24"   23 58'48"
 13:30         20.50          130.19    -21.91   204 11'23"   23 36' 0"
 13:35         20.58          131.44    -21.91   205 22'47"   23 11'59"
 13:40         20.67          132.69    -21.91   206 33' 0"   22 47'23"
 13:45         20.75          133.94    -21.91   207 43'11"   22 21' 0"
 13:50         20.83          135.19    -21.91   208 52'12"   21 53'59"
 13:55         20.92          136.44    -21.91   210  0'35"   21 26'24"

 14: 0         21.00          137.69    -21.91   211  8'23"   20 57'36"
 14: 5         21.08          138.94    -21.91   212 15'35"   20 27'36"
 14:10         21.17          140.19    -21.91   213 22'12"   19 56'24"
 14:15         21.25          141.44    -21.92   214 28'11"   19 24'36"
 14:20         21.33          142.69    -21.92   215 33' 0"   18 52'12"
 14:25         21.42          143.94    -21.92   216 37'12"   18 18'35"
 14:30         21.50          145.19    -21.92   217 40'48"   17 44'23"
 14:35         21.58          146.44    -21.92   218 43'47"   17  8'59"
 14:40         21.67          147.69    -21.92   219 46'12"   16 33' 0"
 14:45         21.75          148.94    -21.92   220 47'23"   15 56'23"
 14:50         21.83          150.19    -21.92   221 48' 0"   15 18'36"
 14:55         21.92          151.44    -21.92   222 48'36"   14 40'11"










20: PropVanes, Site tower, Sat 01-Dec-2001 12:10:21 MST, Prop boom angles
Theodolite readings:

1m  357 deg 29 min 50 sec
2m  357 deg 29 min 10 sec
5m  357 deg 29 min 40 sec
10m 357 deg 29 min 15 sec
15m 357 deg 28 min 40 sec
20m 357 deg 28 min 40 sec

Add 180 and convert to degrees to get boom azimuths:

1m  177.497 deg
2m  177.486 deg
5m  177.494 deg
10m 177.488 deg
15m 177.478 deg
20m 177.478 deg


************************************************
Setting boom angle for 1m prop:

rserial cosmos 204

W:0004    560    73   409   50 00d\n\r
W:0004    575    75   396   50 00W\n\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 502  counts\n\r
     Boom angle  = 0.0 degrees\n\r
     True North  = 502  counts\n\r
     Prop Pitch  = 307  mm/rev\n\r
\n\rd1775\r
D1775\r
\n\r
q\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 502  counts\n\r
     Boom angle  = 177.5 degrees\n\r
     True North  = 250  counts\n\r
     Prop Pitch  = 307  mm/rev\n\r
\n\r

******************************************

2m prop:

rserial cosmos 206

(didn't do Q command before changing boom angle)

W:0001  00589  0077  0363  050  00Y\n\r
W:0001  00596  0078  0365  050  00X\n\r
W:0001  00612  0080  0364  050  00k\n\r
\n\r
d1775\r
D1775\r
\n\r
q\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 148  counts\n\r
     Boom angle  = 177.5 degrees\n\r
     True North  = 2224  counts\n\r
     Prop Pitch  = 306  mm/rev\n\r
    Direction Device = RESOLVER\n\r
\n\r
\r
\r
\n\r
W:0001  00650  0085  2275  050  40]\n\r
W:0001  00634  0083  2219  050  40_\n\r
W:0001  00657  0086  2188  050  40R\n\r

***********************************************
5m prop:

 rserial cosmos 209
*** use ^d or ^c to terminate rserial
send socket open to goldenrod
recv socket open to goldenrod
 86   50 00w\n\r
W:0002    686    90   114   50 00e\n\r
W:0002    648    85   169   50 00Y\n\r
W:0002    648    85   170   50 00a\n\r
W:0002    632    83   157   50 00e\n\r
\n\r
q\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 402  counts\n\r
     Boom angle  = 0.0 degrees\n\r
     True North  = 402  counts\n\r
     Prop Pitch  = 305  mm/rev\n\r
\n\r
d1775\r
D1775\r
\n\r
q\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 402  counts\n\r
     Boom angle  = 177.5 degrees\n\r
     True North  = 150  counts\n\r
     Prop Pitch  = 305  mm/rev\n\r
\n\r
\r
\r
\n\r
W:0002    922   121  1968   50 00\0x7f\n\r
W:0002    953   125  1896   50 00w\n\r
W:0002    937   123  1893   50 00z\n\r
W:0002    922   121  1859   50 00@\n\r


*********************************** 

10m prop
rserial marigold 203
*** use ^d or ^c to terminate rserial
send socket open to arnica
recv socket open to arnica
W:0003    808   105   203   50 00\\n\r
W:0003    800   104   215   50 00b\n\r
W:0003    777   101   246   50 00T\n\r
W:0003    739    96   246   50 00Y\n\r
W:0003    716    93   225   50 00d\n\r
W:0003    693    90   224   50 00d\n\r
W:0003    662    86   220   50 00g\n\r
W:0003    662    86   204   50 00e\n\r
W:0003    662    86   206   50 00c\n\r
W:0003    646    84   219   50 00_\n\r
W:0003    639    83   244   50 00`\n\r
\n\r
q\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 496  counts\n\r
     Boom angle  = 0.0 degrees\n\r
     True North  = 496  counts\n\r
     Prop Pitch  = 308  mm/rev\n\r
\n\r
d 1775\r
D 1775\r
\n\r
q\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 496  counts\n\r
     Boom angle  = 177.5 degrees\n\r
     True North  = 244  counts\n\r
     Prop Pitch  = 308  mm/rev\n\r
\n\r
\r
\r
\n\r
W:0003    708    92  1991   50 00I\n\r
W:0003    669    87  1885   50 00}\n\r
W:0003    639    83  1754   50 00I\n\r
W:0003    654    85  1787   50 00D\n\r


************************************************

15m prop

rserial marigold 205
*** use ^d or ^c to terminate rserial
send socket open to arnica
recv socket open to arnica
71   141   282   50 00I\n\r
W:0006   1071   141   288   50 00C\n\r
W:0006   1056   139   268   50 00{\n\r
W:0006   1056   139   264   50 00\0x7f\n\r
W:0006   1071   141   265   50 00H\n\r
W:0006   1079   142   273   50 00@\n\r
W:0006   1064   140   274   50 00G\n\r
W:0006   1064   140   257   50 00F\n\r

\n\r
q\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 253  counts\n\r
     Boom angle  = 0.0 degrees\n\r
     True North  = 253  counts\n\r
     Prop Pitch  = 304  mm/rev\n\r
\n\r
d1775\r
D1775\r
\n\r
q\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 253  counts\n\r
     Boom angle  = 177.5 degrees\n\r
     True North  = 001  counts\n\r
     Prop Pitch  = 304  mm/rev\n\r
\n\r
\r
\r
\n\r
W:0006    934   123  2060   50 00F\n\r
W:0006    988   130  2044   50 00}\n\r
W:0006   1026   135  2062   50 00x\n\r
W:0006   1026   135  2049   50 00s\n\r
W:0006   1018   134  2033   50 00z\n\r

*****************************************

20m prop

 rserial marigold 207
*** use ^d or ^c to terminate rserial
send socket open to arnica
recv socket open to arnica
W:0007  00826  0108  0311  050  40a\n\r
W:0007  00818  0107  0328  050  40Y\n\r
W:0007  00826  0108  0342  050  00a\n\r
W:0007  00833  0109  0351  050  00b\n\r
W:0007  00826  0108  0359  050  00Y\n\r
W:0007  00810  0106  0357  050  00d\n\r

Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 2943  counts\n\r
     Boom angle  = 0.0 degrees\n\r
     True North  = 2943  counts\n\r
     Prop Pitch  = 306  mm/rev\n\r
    Direction Device = RESOLVER\n\r
\n\r
d1775\r
D1775\r
\n\r
q\r
Q\r
\n\r
Parameters \n\r
     Sample rate = 250 Hz\n\r
     Xmit rate   = 0050 samples\n\r
     Boom align  = 2943  counts\n\r
     Boom angle  = 177.5 degrees\n\r
     True North  = 923  counts\n\r
     Prop Pitch  = 306  mm/rev\n\r
    Direction Device = RESOLVER\n\r
\n\r
\r
\r
\n\r
W:0007  00627  0082  1927  050  40S\n\r
W:0007  00650  0085  1935  050  40U\n\r
W:0007  00665  0087  1948  050  00M\n\r
W:0007  00665  0087  1960  050  40O\n\r













19: Sonics, Site tower, Sat 01-Dec-2001 11:49:05 MST, Tower sonic boom angles
Theodolite readings:
2m  87 deg 51 min 20 sec
10m 87 deg 51 min 15 sec
20m 87 deg 51 min 15 sec

Add 180 and convert to degrees to get boom azimuths:

2m:  267.855 deg
10m: 267.854 deg
20m: 267.854 deg

Add 90 to get Vazimuth: 357.855 deg


Here is the sunangle data used:

		Sun Angle Data for

		earth location....latitude:   40,39,32
				  longitude: 106,19,26
		on date: 12/ 1/2001 

local time,    Z time         Grnwich    earth      local sun position  
hour/minute                  Hour Angle  decl.      azimuth    elev.  
-----------    ------        ----------  -----      ------------------

  9: 0         16.00           62.71    -21.88   138 22'47"   15 28'48"
  9: 5         16.08           63.96    -21.88   139 24' 0"   16  6' 0"
  9:10         16.17           65.21    -21.88   140 25'48"   16 42'36"
  9:15         16.25           66.46    -21.88   141 28'11"   17 18'35"
  9:20         16.33           67.71    -21.88   142 31'12"   17 53'24"
  9:25         16.42           68.96    -21.88   143 34'48"   18 27'36"
  9:30         16.50           70.21    -21.89   144 39'35"   19  0'36"
  9:35         16.58           71.46    -21.89   145 44'24"   19 33'35"
  9:40         16.67           72.71    -21.89   146 50'24"   20  4'47"
  9:45         16.75           73.96    -21.89   147 56'59"   20 35'23"
  9:50         16.83           75.21    -21.89   149  4'48"   21  5'23"
  9:55         16.92           76.46    -21.89   150 12'36"   21 33'35"

 10: 0         17.00           77.71    -21.89   151 20'59"   22  1'48"
 10: 5         17.08           78.96    -21.89   152 30'35"   22 28'11"
 10:10         17.17           80.21    -21.89   153 40'48"   22 53'59"
 10:15         17.25           81.46    -21.89   154 51'36"   23 18'35"
 10:20         17.33           82.71    -21.89   156  3' 0"   23 42'36"
 10:25         17.42           83.96    -21.89   157 15' 0"   24  4'47"
 10:30         17.50           85.21    -21.89   158 27'36"   24 26'24"
 10:35         17.58           86.46    -21.89   159 40'48"   24 46'48"
 10:40         17.67           87.71    -21.89   160 54'35"   25  5'23"
 10:45         17.75           88.96    -21.89   162  8'23"   25 23'24"
 10:50         17.83           90.21    -21.89   163 23'23"   25 40'12"
 10:55         17.92           91.45    -21.89   164 39' 0"   25 56'24"

 11: 0         18.00           92.70    -21.89   165 54'35"   26 10'47"
 11: 5         18.08           93.95    -21.90   167 10'48"   26 23'59"
 11:10         18.17           95.20    -21.90   168 27'36"   26 36' 0"
 11:15         18.25           96.45    -21.90   169 44'24"   26 46'11"
 11:20         18.33           97.70    -21.90   171  1'48"   26 55'47"
 11:25         18.42           98.95    -21.90   172 19'11"   27  4'12"
 11:30         18.50          100.20    -21.90   173 36'36"   27 10'47"
 11:35         18.58          101.45    -21.90   174 54'35"   27 16'48"
 11:40         18.67          102.70    -21.90   176 13'11"   27 21' 0"
 11:45         18.75          103.95    -21.90   177 31'12"   27 23'59"
 11:50         18.83          105.20    -21.90   178 49'48"   27 25'47"
 11:55         18.92          106.45    -21.90   180  7'47"   27 26'24"








18: Adams, Site tower, Sat 01-Dec-2001 10:18:24 MST, Swapped ironics card in marigold
Arnica, the ironics card in marigold died last night.

This morning at about 10:00 MST the spare ironics (SN 422016)
was swapped in.



16: Sonics, Site stn2, Fri 30-Nov-2001 18:04:24 MST, bare stuff
sonic height now 2.97m.  We will change it back down to 2m later.
sonic azimuth 271 (handheld compass, decl=12E); 270 (KVH #3)

15: soil, Site all stns, Sat 05-Aug-2000 08:53:54 PDT, Soil water gravimetric data
FLOSS01 Soil Gravimetric water fraction. 
All the soil moisture data is to be entered in this logbook file


Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.  
 62 SO  06 Dec 01 Bare    0-2   A1 18.8    16.5    2.3    0.14
                                A2 19.6    17.4    2.2    0.13  = 0.14
                          2-5   B1 23.8    20.8    3.0    0.14
                                B2 24.7    21.7    3.0    0.14  = 0.14
                          5-8   C1 34.8    31.7    3.1    0.10
                                C3 36.4    33.7    2.7    0.08  = 0.09
                          8-10+ D1 30.9    28.0    2.9    0.10
                                A3 30.6    28.2    2.4    0.09 	= 0.10 	                                      
---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
 68 SO  08 Dec 01 tower   0-3   B1 11.8     8.1    3.7    0.05
                                B2 12.3     6.6    5.7    0.86  = ?
                          3-6   C1 32.9    29.4    3.5    0.12
                                C3 32.0    28.1    3.9    0.14  = 0.13
                          6-8+  D1 40.8    37.7    3.1    0.08
                                A3 36.1    33.3    2.8    0.08  = 0.08

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
 71 SO  08 Dec 01 tower 0-1.5   A1 13.3     2.9   10.4    3.6
                                A2 10.9     2.4    8.5    3.5   = 3.6
                        1.5-4.5 B1 27.0    12.2   14.8    1.2
                                B2 15.4     9.7    5.7    0.59  = 0.90
                        4.5-7.5 C1 47.9    43.1    4.8    0.11
                                C3 33.8    30.5    3.3    0.11  = 0.11
                        7.5-9.5 D1 33.8    31.7    2.1    0.07
                                A3 34.7    32.8    1.9    0.06  = 0.07

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
118 SS  17 Jan 02 tower   0-5   A1 14.3    15.9
                                A2 16.3    17.4                     ? 
                                A3 16.3    18.0


---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
118 SS  17 Jan 02 sage    0-5   B1 29.7    25.7   4.0     0.16
                                B2 31.0    26.4   4.6     0.17   = 0.16
                                B3 31.3    27.0   4.3     0.16

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
141 SO  18 Feb 02 tower   0-9.5 A1 36.3    33.3    3.0    0.09
                                A2 26.7    24.3    2.4    0.10   = 0.11
                                A3 28.8    25.4    3.4    0.13

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
146 TD  20 Feb 02 sage    0-5   A1 19.4    16.5    2.9    0.18
                                A2 20.4    17.4    3.0    0.17  = 0.18
                                A3 23.4    19.7    3.7    0.19

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
146 TD  20 Feb 02 bare    0-5   B1 19.6    16.7    2.9    0.17     
                                B2 23.9    20.1    3.8    0.19  = 0.18
                                B3 27.1    22.7    4.4    0.19

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
146 TD  20 Feb 02 tower   0-5   C1 17.6    13.2    4.4    0.33
                                C2 17.0    12.7    4.3    0.34  = 0.35
                                C3 16.6    12.0    4.6    0.38

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
147 TD  22 Feb 02 tower   0-2   D1  5.8     3.6    2.2    0.61
                                D2  4.7     2.8    1.9    0.68  = 0.64
                                D3  5.9     3.6    2.3    0.64
                          2-4   E1 14.9    11.9    3.0    0.25
                                E2 17.3    13.7    3.6    0.26  = 0.25
                                E3 19.0    15.2    3.8    0.25
                          4-8   F1 19.2    17.1    2.1    0.12
                                F2 18.7    16.7    2.0    0.12  = 0.12
                                F3 20.4    18.3    2.1    0.11
                          8-16  G1 26.3    24.9    1.4    0.06
                                G2 23.0    21.8    1.2    0.06  = 0.06
                                G3 29.0    27.4    1.6    0.06

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD  27 Feb 02 Sage    0-2   A1 22.1    18.1
                                A2 25.9    21.2
                                A3 26.1    21.5

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD  27 Feb 02 Sage    2-5   B1 23.8    21.5
                                B2 21.9    19.8
                                B3 26.1    20.6

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD 27 Feb 02  Sage    5-10  C1 30.7    27.9
                                C2 29.6    26.8
                                C3 22.9    20.6

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD 27 Feb 02  Bare    0-2   D1 11.3     9.5
                                D2  8.9     7.6
                                D3 11.0     9.3

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD 27 Feb 02  Bare    2-5   E1 21.7    18.6
                                E2 19.9    16.8
                                E3 19.8    16.9

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD 27 Feb 02  Bare    5-10  F1 24.2    21.4
                                F2 24.7    21.8
                                F3 24.5    21.7

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
162 TD  28 Feb    Tower   0-2   A1 3.9      2.0
                                A2 5.1      2.5
                                A3 3.9      1.9

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
162  TD  28 Feb    Tower   2-5   B1 7.0      5.7
                                 B2 9.1      7.6
                                 B3 7.2      6.1

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
162 TD  28 Feb    Tower   5-10  C1 24.5    23.2
                                C2 27.9    26.3
                                C3 27.6    26.0

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
171     13 Mar    Tower   0-2   IA   9.2    3.5
                                IB   9.6    3.6

                          2-5  IIA  24.6   16.9
                               IIB  31.6   23.2
                                
                         5-10 IIIA  40.2   36.7
                              IIIB  39.2   35.9

Oven-dried aliquots weighed on 26 Mar 02. See note in 171
---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
177 TD  26 Mar    Tower   0-2  IA   8.7   4.2
                               IB  10.0   4.6
                               IC  11.0   4.9

                          2-5 IIA  29.3  22.9
                              IIB  35.1  27.2
                              IIC  31.8  23.9

                         5-10 IIIA 30.2  25.2
                              IIIB 36.6  29.5
                              IIIC 34.1  28.1
Oven-dried aliquots weighed on April 2. GDM





















14: Status, Site all stns, Thu 29-Nov-2001 17:49:08 MST, status
Today's work:
- installed all T/RHs on main
- intalled all props on main
- installed 2m CSAT & KH20 on main
- set up rad stand (but not installed in final position)
- set up bare site infrastructure and some sensors

Other:
- Larry Mahrt and Dean Vickers arrived, staked out their sites and organized their equipment
- Sean Burns arrived, will set up tomorrow
- Mark Lanier stopped by for a visit

13: Base, Site base, Thu 29-Nov-2001 14:28:14 MST, hourly.csh, daily.csh, and SSH
Crontab executes these scripts on $ASTER/projects/FLOSS01/scripts:

  hourly.csh: executed at XX:15 every hour from 06:00 to 22:00.
  daily.csh: executed every morning at 06:20 MST

Add commands to these scripts do periodic tasks.

One script that is executed from hourly is update_ip_addr:

  update_ip_addr:  does an ssh to maclean@linus and copies the current
		IP address to /net/aster/projects/FLOSS01/floss_ip_addr

   

12: Communications, Site base, Thu 29-Nov-2001 13:53:20 MST, ISP and LAN-modem
The 3com LAN-modem has been configured to dial-up the Earthlink
ISP in Steamboat Springs.

To configure the modem from a system connected to it,
go to this URL with a browser:

  http://3com.oc.lanmodem

I've added it as "Lan-modem" to aster's personal toolbar in netscape, so
you can just click on that.

The administrator password is the same as the root password on aster.

ISP info:

  steamboat phone number: 871-2640
  username: see gordon
  password: see gordon

We've set the call duration to 15 minutes.



11: Adams, Site tower, Thu 29-Nov-2001 03:47:18 GMT, Ironics failures & light bulb heaters
The Ironics cards, both arnica and goldenrod, have been dying.  They
stay up for a while, maybe 15 minutes or 2 hours, and then die.

Once goldenrod rebooted itself, but the adam code does not support
an ironics-only reboot, and so the serial data will not
get through.

One change from the testing at FLAB is the drop in the
internal adam temperature.

This afternoon we added some serial sensors to cosmos, so now
we can detect in the netcdf file if the ironics is running.

This evening after sunset Tbox.cosmos was dropping from 5.9 C.
At 3.6 C the ironics rebooted.

At 11/28 8:35 pm MST we placed a 40Watt bulb in the VME cage, between
the Datel and the filter board. Overnight Tbox.cosmos got up to
24C, with an outside temp of around -10C.  The Ironics stayed up
all night!

At 10:00 am MST, 11/29, replaced the 40W bulb with a 7.5 W, and put
a 7.5W buld in marigold. were placed in both adams.




10: Status, Site tower, Thu 29-Nov-2001 02:25:48 GMT, Tower heights
I crudely set heights of booms today.  Here is what I did (NOTE THE FUTURE WORK LISTED IN 7-9):
1. The wood beams for the tower base are 10cm high
2. The metal plates under the jacks are ~1.5cm thick (probably 1/2")
3. The height of the top of the tube at the top of the first ladder above the metal plates for each of the 4 legs were: 199, 196, 197, & 202 cm
4. I average the values in 3. and add to the values of 1. and 2. to get a nominal height above ground of the top of the first ladder of 210 cm
5. We measured the height of one section as 184cm
6. Thus, I get the following heights above ground of the top of the tube at the top of each of the ladder sections (in cm):
210
394
578
762
946
1120
1304
1488
1672
1856
2040
2224
2408
These are labeled using duct tape on the tubes.
7. We then attempted to position each of the sensors as close as possible to the specified sensor heights, 0.5, 1, 2, 5, 10, 15, & 20m.  For the sensors up to 2m, we got to within 10cm by special clamping (still not perfect due to conflicts with brackets, etc.  For the sensors above 2m, we should be within 50cm.  We assumed that the sonics were 5cm above the boom, the props 50cm above the boom, and the T/RHs 50cm below the boom.  Also, anemomber yokes make the boom height ~6cm lower than the tube it is clamped to and T/RH yokes (vertical) set the boom ~7cm from the bottom (or top) edge.  Note that CSATs end up with the array almost perfectly aligned with the center of the tube that the yoke is clamped to.  All these relative heights should be checked as well.
8. I'll have to measure the exact positions (still relative to the duct-tape labels) later.
9. Overall, I am reasonably comfortable with this approach.  However, the 2m sonic will end up closer to 2.5m due to the tube available to clamp to.  Also, the terrain to the south and west slope up a bit, so the sensors at the end of 3 and 4m long booms may be 10-15cm closer to the local surface.  I'll try to determine this as well at a later time.


9: Status, Site tower, Wed 28-Nov-2001 23:56:08 GMT, status
Stuff done Monday:
trailer didn't get here until 4:30.  powered up and porch set-up by 5:45.
(overcast, flurries, cold)

Stuff done Tuesday:
- power to tower
- phone connected
- base computers installed
- adams up and transmitting Tbox by ~5:00
(overcast, flurries, cold)

Stuff done today:
- hung all booms on tower
- connected T/RH at 0.5 and 1.0m
- rad stand erected near tower base, rads attached
John and Tony show up ~2:00
(clear, very cold -- -30 last night -- wind started up about 12:30)

8: Adams, Site all stns, Tue 20-Nov-2001 16:15:06 GMT, manual control note
The ttl manual control is reverse logic.
So ttl_on turns power off and ttl_off turns power on.
I confirmed this on marigold. I assume it is the same
for cosmos.
7: Adams, Site tower, Thu 15-Nov-2001 06:06:52 GMT, manual control of TTL ports
Three programs provide for manual control of the ttl ports on an adam.
The manual settings override what ever ttl program is running
on the adam (as specified in channel_config).

To return a ttl port to the control of the ttl program, use ttl_reset.

  ttl_on adam ports...    # Turns on one or more ports, numbered from 0-2
  ttl_off adam ports...
  ttl_reset adam ports	   # returns one or more ports to the normal
			     control of the ttl program.

Example:

  ttl_on marigold 0 2	# turn on ports 0 and 2
  ttl_off marigold 0	# turn off port 0
  ttl_reset marigold 2	# put port 2 back under control of the adam
6: Power, Site base, Thu 15-Nov-2001 05:59:36 GMT, UPS software
The black cable with DB9 connectors should be connected between
the UPS and the serial port connection on the back of the dell.

The UPS monitor software runs automatically on bootup.

I have tested it with the Smart-UPS 620.  I don't know if the
software needs to be reconfigured to use with the bigger UPS.

The software and log files are in /usr/lib/powerchute.

For display, run: /usr/lib/powerchute/xpowerchute

To reconfigure (if you switch to a bigger ups): Config.sh

The log files are:
	powerchute.log	# power events
	powerchute.dat  # time series of voltage, etc - potentially useful!


5: Communications, Site all stns, Sat 10-Nov-2001 23:44:39 GMT, Breezecom settings
AP: Access point, this will be the radio at the base, in the trailer
SA: Station adaptor, this is the radio out at the tower.

Community strings:
  read: ncar/atds
  write: (see gordon)

  WLAN parameters -> ESS ID: see gordon
  WLAN parameters -> fixed channel, #11

For testing in the building, I have 2 little dipole antennals
on each radio:
  WLAN parameters -> Transmit Diversity -> both antennas

Security:
  Default key ID: First
  WEP Keys: see gordon
    1st: 
    2nd: 
    3rd: 
    4th: 


Breezecoms were showing a high FCS error rate.  According to the
DS11 manual this is a CRC error statistic.  The adams seemed to
be suffering, losing data, and crashing.

Tried various things, which didn't work, including 1/2 meter dipole antenna
on AP, moving radios further apart, resetting to factory defaults.

What did help was changing the channel number from 6 to 11.  It appears
that another wireless ethernet in the building is using channel 6.

According to the breezecom manual (Appendix, page B-2), only channels
1,6 and 11 are non-overlapping.

Settings:
  SA: WLAN Parameters -> Channel Settings -> Fixed Channel, #11
  AP: WLAN Parameters -> Channel 11

To view the FCS statistic, select the SA, click on the "Counters" tab.
A useful tool is to click on the FCS number and drag it to the
graph image.  Once the graph appears, click on scale to change
the scale to 0-10 counts. Currently we're seeing an FCS error 
rate of 5%, with one or two errors every 5 seconds.  This 
seems to be acceptable.


4: Communications, Site all stns, Sat 10-Nov-2001 22:44:47 GMT, Adam/network testing

The adams have been transmitting via the breezecoms since yesterday
(fri Nov 9) afternoon.  The security was set to "secured key",
which SCD recommends.

The network has not been working well since then.
The adams would start reporting lost samples
and crashed several times.

The radios are set up with the indoor dipole antennas.  Until now
I had both antennas connected.   The DS11 program was showing a bunch
of errors.  The FCS Error rate was about 75%.

According to the DS11 manual, only antenna #1 should be connected.
After making that change, things have improved. We'll see how it goes.
The breezecoms are about 8 feet apart in the staging area.

Current error status shown on DS11 program, 

  SA showing 80-85 % signal quality,  RSSI -27dBm (62)
     FCS Error 11%

  AP: FCS Error 11%

It looks like the FCS error rate is a critical number.


3: Adams, Site base, Sat 10-Nov-2001 21:59:05 GMT, Adam VME serial numbers
cosmos:
  backplane #3
  matrix: SN 9013
  ironics iv3234:  SN129043
  datel: 320689

marigold
  backplane: #4
  matrix: SN 940211101
  ironics iv3234: SN 821003
  datel: SN 320589


On 12/1 the Ironics in marigold was replaced with SN: 422016
(which is the card we got from MIT).

2: Base, Site base, Fri 09-Nov-2001 08:15:22 MST, Archive size and scheme
With the 3 sonics on the adams at 60hz the archive sizes are

	marigold 67 Mbyte/8 hr = 201 Mbyte/day
	cosmos   40 Mbyte/8 hr = 120 Mbyte /day

(no thermocouples on cosmos, and prop vanes at 1 Hz)

SteveO says the thermocouples will run at about 5 Hz and would like
to increase the prop vanes to 5 Hz.

12 thermocouples at 5 hz would be additional 23 Mbyte/day.

The additional prop data is about 25 Mbyte/day.

Eve, one campbell at 20 hz (no krypton/heater yet):
	stn2	8.8 Mbyte/ 7.3 hr = 30 Mbyte/day

So the archive will be about 400 Mbyte/day.

This does not include camera images or more than 12 thermocouples.

We will have about 11 Gbyte available on the base system HD,
and 18 Gbyte on the removable HD.

11 Gbyte gives us about 20-25 days archive.  So we can keep a second
copy of the files on the removeable HD, and about every 2-3 weeks
someone could swap in another removeable HD, and bring the
other back to Boulder. Hopefully we wouldn't delete any files
on the base until they have been copied to the mass store.

To do this we will need to buy an another 18 Gb HD and its
plastic holder for the Dell latitude module bay.


1: LOG, Site none, Fri 09-Nov-2001 08:10:41 MST, Logbook created
initial log file /net/aster/projects/FLOSS01/logbook/tklog.log created by USER maclean